### v1

# Covariate balancing estimation and model selection for difference-in-differences approach

Abstract
In causal inference, remarkable progress has been made in difference-in-differences (DID) approaches to estimate the average effect of treatment on the treated (ATT). Of these, the semiparametric DID (SDID) approach incorporates a propensity score analysis into the DID setup. Supposing that the ATT is a function of covariates, we estimate it by weighting the inverse of the propensity score. As one method to make the estimation robust to the propensity score modeling, we incorporate covariate balancing. Then, by attentively constructing the moment conditions used in the covariate balancing, we show that the proposed estimator has doubly robustness. In addition to the estimation, model selection is also addressed. In practice, covariate selection is an essential task in statistical analysis, but even in the basic setting of the SDID approach, there are no reasonable information criteria. Therefore, we derive a model selection criterion as an asymptotically bias-corrected estimator of risk based on the loss function used in the SDID estimation. As a result, we show that a penalty term is derived that is considerably different from almost twice the number of parameters that often appears in AIC-type information criteria. Numerical experiments show that the proposed method estimates the ATT robustly compared to the method using propensity scores given by the maximum likelihood estimation (MLE), and that the proposed criterion clearly reduces the risk targeted in the SDID approach compared to the intuitive generalization of the existing information criterion. In addition, real data analysis confirms that there is a large difference between the results of the proposed method and the existing method.
Keywords: Covariate balancing, Difference-in-differences, Information criterion, Model selection, Propensity score, Semiparametric approach, Statistical asymptotic theory
1 Introduction
The difference-in-differences (DID) approach is one of the most common methods used in fields such as economics and epidemiology to assess the effect of a particular intervention or treatment. In the basic setting, there are two groups: a treatment group receiving a specific treatment and a control group receiving no treatment or other treatment. The pre and post-treatment outcomes of each group are measured and the differences in the outcomes are compared between the groups to estimate the average treatment effect on the treated (ATT). Under the parallel trend assumption, simple estimation is valid. More recently, Abadie (2005) dealt with inference under more naturalized assumptions using information from the covariates, specifically the parallel trend assumption conditioned on the covariates. Then, as a method that does not necessarily require modeling of the covariates to the outcome, they propose the semiparametric DID (SDID) method, which estimates ATT by weighting the inverse of the propensity score.
With or without conditionality, the parallel trend assumption is a key assumption in the DID. This is the assumption that the mean outcomes of a treatment group and a control group would have been similar over time if there had been no treatment. On the other hand, especially in epidemiology, the ignorable treatment assignment assumption has usually been invoked in causal inference (e.g., Robins et al. 1994 and Li et al. 2018). This is the assumption that the potential outcome variable and the treatment assignment variable, which represents the received treatment, become independent conditioned on the covariates. Although neither assumption is verifiable, the validity of the parallel trend assumption can be examined by visual inspection when we have observations over many time points. And the parallel trend assumption naturally holds if it can be assumed that there is no change over time in the expected value of the outcome without the treatment. Needless to say, the parallel trend assumption is considerably weaker than the ignorable treatment assignment assumption, and estimating the ATT under such desirable conditions is the advantage that the DID gained in exchange for collecting pre- and post-treatment outcomes.
The methodological development of the DID has been remarkable in recent years because of such advantages and the increasing amount of available observational study data (see Roth et al. 2023 for a comprehensive review of the DID in recent years). However, for example, Abadie (2005) assumes that the model for the propensity score is correctly specified, but if the model is incorrect, the resulting estimator has a bias. As an extension of Abadie (2005), Sant’Anna and Zhao (2020) assumes not only a regression model for the treatment assignment variable used to obtain the propensity score, but also a regression model for the outcome, and proposes the estimator for the ATT that combines the results of each estimator. Their proposed estimator is consistent if either a propensity score or outcome regression working model is correctly specified, that is, the estimators have doubly robustness.
In this article, as with Abadie (2005), we assume that ATT is a function of covariates and address the semiparametric estimation problem for that function. We assume two regression models as above and give double robustness to the estimation, but unlike the Sant’Anna and Zhao (2020), we incorporate covariate balancing (Imai and Ratkovic 2014), a method that uses propensity scores such that a certain moment condition is satisfied. This method actually estimates the model for the treatment assignment, while assuming, but not estimating, the model for the outcome. Nevertheless, the SDID estimator derived by weighting the inverse of the propensity score, which is obtained by the covariate balancing, will be consistent if either one of the regression models is correctly specified. One point to note is that, in order to satisfy this favorable asymptotic property in DID, the moment condition is constructed in a different way from the usual one. Incidentally, incorporating covariate balancing leads to the estimation of causal effects via weights that always balance the moments of the covariates across groups, which is also preferable from an interpretive point of view (Imai and Ratkovic 2014). We call the proposed methodology the covariate balancing for DID (CBD) method.
Obviously, model selection is an essential task in the DID as well, in a series of data analysis tasks. For example, in estimating the effect of union membership on wages, Houngbedji (2016) treats a linear model with workers’ demographics for causal effects, and explores factors affecting the causal effect on the basis of the values of the regression coefficients and its $p$-value. In other words, the issue of covariate selection is addressed. However, for the information criterion that is the basic tool for model selection, there are no reasonable ones, even for the basic setting of the SDID. Thus, we derive the information criterion for the basic SDID and CBD methods as an asymptotically unbiased estimator of the risk based on the loss function used in the SDID estimation. Note that traditional methods such as GIC (Konishi and Kitagawa 1996) cannot be applied because the loss function includes weights based on the propensity score, which is a random variable. Since the SDID method places importance on being based on the conditional parallel trend assumption, it does not rely on the ignorable treatment assignment condition when deriving the information criterion. This causes some difficulties. As a result, we derive penalty terms that are significantly different from twice the number of parameters, which often appear in AIC-type information criteria including GIC.
In summary, the aims of this article are to develop a new doubly robust estimation method with the advantage of covariate balancing in the SDID method, and to derive model selection criteria for the SDID method, for which no reasonable model selection criteria exist yet. The organization of this article is as follows. First, in section 2, the model and estimation target are specified, and the SDID method proposed in Abadie (2005) is introduced. In section 3, we propose the CBD method and give a theorem showing its doubly robustness. Numerical experiments show that the estimation is robust against model misspecification. In section 4, when the propensity score is known, we derive the model selection criterion as a bias-corrected estimator of the risk that is naturally defined from the loss function used in the SDID method. Then, we also give a theorem that suggests that the penalty term of the model selection criterion is asymptotically valid. As a comparison, we derive the model selection criterion the QICw, which naturally extends the idea of Platt et al. (2013) to the SDID method. Furthermore, through numerical experiments, we verify the approximation accuracy of the penalty term of the proposed criterion from the perspective of estimating the risk, and also verify its performance compared to the QICw from the viewpoint of risk minimization. Section 5 summarizes the results for the CBD method in the same manner as Section 4. In other words, model selection criteria are derived, theorems are given to support their validity, and their performance is evaluated through numerical experiments. Section 6 reports the results of estimating causal effects and model selection for the real data in LaLonde (1986). Finally, Section 7 discusses the future development of the proposed method.
2 Preparation
2.1 Model and Assumptions
Suppose that there are two groups, a treatment group and a control group, and that outcomes are collected before and after the possible treatment time point, respectively. We treat the counterfactual setting, where $y^{[1]}(t)$ denotes the outcome of the treatment group, $y^{[0]}(t)$ the outcome of the control group, $t=0$ the outcome before treatment, and $t=1$ the outcome after treatment. In practice, only one of the outcomes of the treatment or control group can be obtained. Preparing treatment assignment variables such that $d^{[1]}=1\ (d^{[0]}=0)$ if assigned to the treatment group and $d^{[0]}=1\ (d^{[1]}=0)$ if assigned to the control group, the observed outcome collected at each time point can be expressed as
$y(t)=d^{[1]}y^{[1]}(t)+d^{[0]}y^{[0]}(t).$
Letting the change in the outcome variable at two time points in the treatment group be $\Delta^{[1]}=y^{[1]}(1)-y^{[1]}(0)$ and the change in the control group be $\Delta^{[0]}=y^{[0]}(1)-y^{[0]}(0)$, the observed change can be expressed as
$\Delta=d^{[1]}\Delta^{[1]}+d^{[0]}\Delta^{[0]}.$
Here, for the ATT conditioned on the covariate $\bm{x}\ (\in\mathbb{R}^{p})$,
$a(\bm{x})\equiv{\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1],$ (1)
we suppose $g(\bm{x};\bm{\theta})$ $(\bm{\theta}\in\bm{\Theta}\subset\mathbb{R}^{p})$ as a working model to approximate it, and for simplicity, we suppose the working model to be a linear model $\bm{x}^{{\rm T}}\bm{\theta}$. Then, we estimate $\bm{\theta}$ and give the ATT by taking the expectation with respect to $\bm{x}$ in the treatment group.
Now, we give the optimal value of $\bm{\theta}$ as
$\bm{\theta}^{*}\equiv\mathop{\rm argmin}_{\bm{\theta}}{\rm E}[\{{\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]-\bm{x}^{{\rm T}}\bm{\theta}\}^{2}\mid d^{[1]}=1].$ (2)
Abadie (2005) shows that $\bm{\theta}^{*}$ can be consistently estimated by the SDID method using propensity scores $e^{[1]}(\bm{x})\equiv{\rm P}(d^{[1]}=1\mid\bm{x})$ and $e^{[0]}(\bm{x})\equiv{\rm P}(d^{[0]}=1\mid\bm{x})$ under several assumptions, and thus the ATT can be consistently estimated. The assumptions are as follows.
Assumption 1 (Conditional parallel trend).
${\rm E}[y^{[0]}(1)-y^{[0]}(0)\mid\bm{x},d^{[1]}=1]={\rm E}[y^{[0]}(1)-y^{[0]}(0)\mid\bm{x},d^{[0]}=1],$
that is, for the expected value conditioned on $\bm{x}$ of the change in outcome over time without the treatment, the values for the treatment and control groups are the same.
Assumption 2 (Common shocks).
Events occurring between the time points of outcome collection before and after the treatment have the same impact on outcomes in both the treatment group and the control group.
Assumption 3 (Positivity).
$0<{\rm P}(d^{[1]}=1\mid\bm{x})<1,$
that is, the propensity score is not 00 or $1$.
Abadie (2005) showed that using the weights $\rho(\bm{d},\bm{x})\equiv d^{[1]}/e^{[1]}(\bm{x})-d^{[0]}/e^{[0]}(\bm{x})$, then
${\rm E}[\rho(\bm{d},\bm{x})\Delta\mid\bm{x}]$ $={\rm E}[\Delta^{[1]}\mid\bm{x},d^{[1]}=1]-{\rm E}[\Delta^{[0]}\mid\bm{x},d^{[0]}=1]$
$={\rm E}[\Delta^{[1]}-\Delta^{[0]}\mid\bm{x},d^{[1]}=1]$
$={\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]$
$=a(\bm{x})$ (3)
holds. Assumption 1 is used in the second equation. This equation means that by multiplying weights based on the propensity scores to the observed values of change, the conditional ATT can be estimated via DID. Then, from (2) and (3), it is shown that the optimal solution of $\bm{\theta}$ can be expressed as
$\bm{\theta}^{*}=\mathop{\rm argmin}_{\bm{\theta}}{\rm E}[e^{[1]}(\bm{x})\{\rho(\bm{d},\bm{x})\Delta-\bm{x}^{{\rm T}}\bm{\theta}\}^{2}].$ (4)
From this equation,
${\rm E}[e^{[1]}(\bm{x})\bm{x}\{\rho(\bm{d},\bm{x})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$ $={\rm E}[e^{[1]}(\bm{x})\bm{x}\{{\rm E}[\rho(\bm{d},\bm{x})\Delta\mid\bm{x}]-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\{a(\bm{x})-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
$=\bm{0}$ (5)
is satisfied, and therefore, the equation
${\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]\bm{\theta}^{*}$ (6)
is obtained.
2.2 Semiparametric estimation
Given an i.i.d. sample of size $n$, the $i$-th sample is denoted with the subscript $i$. Abadie (2005) proposed
$\hat{\bm{\theta}}=\Bigg{\{}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\Bigg{\}}^{-1}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}$ (7)
as a natural estimator of $\bm{\theta}^{*}$ by solving (4). When the weights based on the propensity scores are multiplied to the differences of the observed values, a term corresponding to the difference in difference appears, thus this method is called the Semiparametric DID (SDID) estimator. The following regular conditions are assumed.
Assumption 4.
$(\mathrm{i})$ $\bm{\theta}^{*}$ is interior of the compact set $\bm{\Theta}\subset\mathbb{R}^{r}$. $(\mathrm{ii})$ ${\rm E}[y(t)^{2}]<\infty$ for all $t$. $(\mathrm{iii})$ ${\rm E}[\bm{x}]$ is bounded and ${\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]$ is nonsingular.
From Assumption 4 and (6),
$\hat{\bm{\theta}}=\Bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\Bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}$ $\stackrel{{\scriptstyle p}}{{\to}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\rho(\bm{d},\bm{x})\Delta]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]\bm{\theta}^{*}$
$=\bm{\theta}^{*}$
is obtained, and therefore $\hat{\bm{\theta}}$ is a consistent estimator of $\bm{\theta}^{*}$. In practice, the propensity score is often unknown. In such cases, the propensity score is modeled as $e^{[1]}(\bm{x};\bm{\alpha})={\rm logit}(\bm{x}^{{\rm T}}\bm{\alpha})$, the parameter $\bm{\alpha}$ is estimated as $\hat{\bm{\alpha}}$ by the maximum likelihood estimation (MLE), and the propensity score in (7) is replaced by $e^{[1]}(\bm{x};\hat{\bm{\alpha}})$ to obtain a semiparametric DID estimator. If the propensity score model is correctly specified, i.e., it can be written as
${\rm P}(d^{[1]}=1\mid\bm{x})={\rm logit}(\bm{x}^{{\rm T}}\bm{\alpha}^{*}),$ (8)
and if $\hat{\bm{\alpha}}$ converges to $\bm{\alpha}^{*}$ with some further assumptions, then $\hat{\bm{\theta}}$ is still a consistent estimator of $\bm{\theta}^{*}$.
3 Covariate balancing for difference-in-difference approach
3.1 Estimation method and statistical properties
Propensity scores are often estimated by the MLE, but if the model for the propensity score is misspecified, the estimator of the ATT has bias. Therefore, we deal with robust methods for the model misspecification of the propensity score. Since doubly robust estimation incorporating estimators of outcome regressions has been addressed in Sant’Anna and Zhao (2020) and Ning et al. (2020a), here we construct doubly robust estimators by incorporating covariate balancing (Imai and Ratkovic 2014). Let $\bm{{\rm O}}$ be a zero matrix and $\bm{d}=(d^{[0]},d^{[1]})$, and consider the following moment conditions:
$\bm{H}^{[1]}(\bm{d},\bm{x};\bm{\alpha})\equiv{\rm E}\bigg{[}e^{[1]}(\bm{x};\bm{\alpha})\bigg{\{}\frac{d^{[1]}}{e^{[1]}(\bm{x};\bm{\alpha})}-1\bigg{\}}\bm{x}\bm{x}^{{\rm T}}\bigg{]}=\bm{{\rm O}},$
$\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})\equiv{\rm E}\bigg{[}e^{[1]}(\bm{x};\bm{\alpha})\bigg{\{}\frac{d^{[0]}}{e^{[0]}(\bm{x};\bm{\alpha})}-1\bigg{\}}\bm{x}\bm{x}^{{\rm T}}\bigg{]}=\bm{{\rm O}}.$
From these equations, ${\rm E}[d^{[1]}\bm{x}\bm{x}^{{\rm T}}]={\rm E}[\{e^{[1]}(\bm{x};\bm{\alpha})/e^{[0]}(\bm{x};\bm{\alpha})\}d^{[0]}\bm{x}\bm{x}^{{\rm T}}]$ is obtained, which means that the expected value of $\bm{x}\bm{x}^{{\rm T}}$ in the treatment group is adjusted to match the expected value of $\bm{x}\bm{x}^{{\rm T}}$ in the control group through weighting by the propensity score. In the original covariate balancing (Imai and Ratkovic 2014), the first-order moments of the covariates are balanced, that is, the $\bm{x}\bm{x}^{{\rm T}}$ in the expectation is naturally set to $\bm{x}$. In this article, the second-order moment of the covariates is balanced.
The semi-vectorization of the upper triangular part of the matrix is written as ${\rm vech}(\cdot)$. Since $\bm{x}\bm{x}^{{\rm T}}$ is a symmetric matrix and defining
$\bm{h}(\bm{d},\bm{x};\bm{\alpha})\equiv({\rm vech}\{\bm{H}^{[1]}(\bm{d},\bm{x};\bm{\alpha})\}^{{\rm T}},{\rm vech}\{\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})\}^{{\rm T}})^{{\rm T}},$
the moment condition can be written as
${\rm E}\{\bm{h}(\bm{d},\bm{x};\bm{\alpha})\}=\bm{0}.$ (9)
[…]
$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})=\bm{0},$ (10)
[…]
$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})\approx\bm{0}$ (11)
is used. Specifically, $\hat{\bm{\alpha}}$ can be obtained by the generalized method of moments (Hansen 1982) or the generalized empirical likelihood method (Qin and Lawless 1994).
Here, we use the generalized method of moments and propose $\hat{\bm{\alpha}}^{\rm CB}\equiv\mathop{\rm argmin}_{\bm{\alpha}}\{\bm{h}_{n}(\bm{\alpha})^{{\rm T}}\bm{W}_{n}\bm{h}_{n}(\bm{\alpha})\}$ as the estimator of $\bm{\alpha}$ by the Covariate balancing (CB) method, where $\bm{h}_{n}(\bm{\alpha})\equiv 1/n\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})$. Defining $\bm{G}_{n}(\bm{\alpha})\equiv\partial\bm{h}_{n}(\bm{\alpha})/\partial\bm{\alpha}$, $\hat{\bm{\alpha}}^{\rm CB}$ satisfies
$\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{h}_{n}(\hat{\bm{\alpha}}^{\rm CB})=\bm{0}.$ (12)
Here, $\bm{W}_{n}$ is an arbitrary semi-positive definite matrix, the simplest of which is the identity matrix, but the optimal matrix, $\{1/n\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})^{{\rm T}}\}^{-1}$, provides the most efficient estimation of $\bm{\alpha}$. However, this optimal matrix is not always practical because it may be too close to the degenerate matrix depending on the data, causing instability in estimation.
Remark 1.
The idea of using constraints that approximately satisfy the empirical version of the moment condition to estimate the parameters of the propensity score has also been used in Ning et al. (2020b). In that article, dealing with the case of high dimensional covariates, they propose a doubly robust estimation method that combines the semiparametric estimation method of ATT with the sparse estimation method. It is also assumed that the outcome is expressed in a linear model of covariates, and to improve the convergence speed of the Horvitz-Thompson estimator, the weak covariate balancing property:
$\frac{1}{n}\sum_{i=1}^{n}\biggl{(}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-1\biggr{)}\bm{\alpha}^{{\rm T}}\bm{x}_{i}\approx 0$
is included as a constraint in estimating the parameter of the propensity score.
Here, we state the regularity conditions of the generalized method of moments (Newey and McFadden 1994), which are necessary to describe the properties of the proposed estimation method.
Assumption 5 (Consistency).
$(\mathrm{i})$ The $\bm{W}_{n}$ converges to $\bm{W}$, which is a semi-positive definite matrix, in probability. Also, $\bm{\alpha}=\bm{\alpha}^{\dagger}$ if and only if ${\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})]=\bm{0}$. $(\mathrm{ii})$ The $\bm{\alpha}$ is the interior of the compact set $\bm{A}\subset\mathbb{R}^{p}$. $(\mathrm{iii})$ The $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is differentiable with probability 1 for each $\bm{\alpha}\in\bm{A}$, that is, $e^{[1]}(\bm{x};\bm{\alpha})$ is differentiable with probability 1. $(\mathrm{iv})$ ${\rm E}[\sup_{\bm{\alpha}\in\bm{A}}\|\bm{h}(\bm{d},\bm{x};\bm{\alpha})\|]<\infty$.
Assumption 6 (Asymptotic normality).
$(\mathrm{i})$ $\bm{\alpha}^{\dagger}$ is interior of $\bm{A}$. $(\mathrm{ii})$ $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is differentiable in a neighborhood $\Gamma$ of $\bm{\alpha}^{\dagger}$, with probability approaching one. $(\mathrm{iii})$ ${\rm E}[\|\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\|^{2}]<\infty$. $(\mathrm{iv})$ ${\rm E}[\sup_{\bm{\alpha}\in\Gamma}\|\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}\|_{\rm F}]<\infty$. $(\mathrm{v})$ $\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\allowbreak\bm{G}(\bm{\alpha}^{\dagger})$ is nonsingular for $\bm{G}(\bm{\alpha})\equiv{\rm E}[\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}]$.
Under the above regularity conditions, the proposed method has the following double robustness.
Theorem 1.
Replacing the propensity score in (7) by $e^{[1]}(\bm{x};\hat{\bm{\alpha}}^{\rm CB})$, we name the covariate balancing for DID (CBD) estimator and denote it by $\hat{\bm{\theta}}^{\rm CBD}$. If the propensity score model $e^{[1]}(\bm{x};\bm{\alpha})$ is correctly specified, or if the change in the outcome follows a linear model of covariates, that is, if
${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\bm{x}^{{\rm T}}\bm{\beta}^{[k]*}$ (13)
holds, then under Assumptions 1-6, this $\hat{\bm{\theta}}^{\rm CBD}$ will converge to $\bm{\theta}^{*}$ in probability. Namely, proposed estimator has double robustness.
Proof.
From Assumption 5, $\hat{\bm{\alpha}}^{\rm CB}$ is consistent with $\bm{\alpha}^{\dagger}$. By expanding $\bm{h}_{n}(\hat{\bm{\alpha}})$ around $\bm{\alpha}^{\dagger}$ and applying Assumption 6 and (12), it is shown to be
$\sqrt{n}(\hat{\bm{\alpha}}^{\rm CB}-\bm{\alpha}^{\dagger})$
=$=-\{\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{G}_{n}(\hat{\bm{\alpha}}^{\rm CB})\}^{-1}\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\sqrt{n}\bm{h}_{n}(\bm{\alpha}^{\dagger})+{\rm o}_{\rm P}(1)$ (14)
$\stackrel{{\scriptstyle\rm d}}{{\to}}{\rm N}(\bm{0},\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\hat{\bm{\alpha}}^{\dagger})\}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{\Omega}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\hat{\bm{\alpha}}^{\dagger})\}^{-1}),$
where $\bm{\Omega}(\bm{\alpha})\equiv{\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})\bm{h}^{{\rm T}}(\bm{d},\bm{x};\bm{\alpha})]$. From the consistency of $\hat{\bm{\alpha}}^{\rm CB}$,
$\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$
=$=\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}$
=$=\bigg{\{}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg{\}}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}+{\rm o}_{\rm P}(1)$ (15)
is also obtained. If the propensity score model is correct and $\bm{\alpha}^{\dagger}=\bm{\alpha}^{*}$ is satisfied, that is, if $e^{[1]}(\bm{x})=e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})$ and $e^{[0]}(\bm{x})=e^{[0]}(\bm{x};\bm{\alpha}^{\dagger})$ hold, then from (5), $\hat{\bm{\theta}}^{\rm CBD}$ is a consistent estimator for $\bm{\theta}^{*}$. On the other hand, even if the model for the propensity score is not correct, when the change in outcome follows a linear model of covariates, that is, when (7) holds,
$a(\bm{x})={\rm E}[\Delta^{[1]}\mid\bm{x},d^{[1]}=1]-{\rm E}[\Delta^{[0]}\mid\bm{x},d^{[0]}=1]=\bm{x}^{{\rm T}}(\bm{\beta}^{[1]*}-\bm{\beta}^{[0]*})$
is obtained from (3). Then, from (6),
$\bm{\theta}^{*}={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]=\bm{\beta}^{[1]*}-\bm{\beta}^{[0]*}$
holds. Therefore,
${\rm E}[e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\bm{x}\{\rho(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
$={\rm E}\biggl{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bigg{\{}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{\rm E}[\Delta_{i}^{[1]}\mid\bm{x}_{i},d^{[1]}_{i}=1]-\frac{d^{[0]}_{i}}{e^{[0]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{\rm E}[\Delta_{i}^{[0]}\mid\bm{x}_{i},d^{[0]}_{i}=1]-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\bigg{\}}\biggr{]}$
$={\rm E}\biggl{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\biggl{\{}\biggl{(}\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\biggr{)}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{1}-\biggl{(}\frac{d^{[0]}_{i}}{e^{[0]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\biggr{)}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{0}\biggr{\}}\biggr{]}$
$={\rm E}\biggl{[}\bm{H}_{1}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{1}-\bm{H}_{0}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{0}\biggr{]}$
$=\bm{0}$
is obtained. Here, the last equality holds from the moment condition (9). From the above, when the model of the propensity score is correctly specified, or when the change in outcomes follows a linear model of covariates,
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}={\rm O}_{\rm P}{(\sqrt{n})}$
holds in (15) and $\hat{\bm{\theta}}^{\rm CBD}$ is consistent. ∎
Remark 2.
Here, we discuss the relationship between the proposed method and recent methods. Sant’Anna and Zhao (2020) proposed a doubly robust estimating method in the study of estimating ATT by SDID, in which the propensity score and outcome are estimated by parametric models, respectively. If one of the models is correctly specified, their proposed estimator is consistent with the ATT. On the other hand, our approach does not estimate the parameters of the model for the outcomes, although the estimator is doubly robust because of the use of covariate balancing. In addition, the second-order moment of the covariates is balanced across groups targeting the treatment group population.
3.2 Numerical experiments
The robustness of the proposed estimating method against model misspecification is examined by simulation. The settings for data generation follow
$x_{1},x_{2}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+\alpha^{*}x_{2})),$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+\beta^{*}x_{1}+\epsilon^{[1]}.$
Throughout this article, all random numbers are generated independently. The parameter $\beta^{*}$ represents the magnitude of the contribution of the covariate $x_{1}$ to the causal effect. The propensity score is estimated using only $x_{1}$, that is, the parameter $\alpha^{*}$ represents the magnitude of the contribution of $x_{2}$ to the true propensity score, which is not included in the model for estimating the propensity score. Note that when $\alpha^{*}\neq 0$, we are considering a model misspecification in which $x_{2}$ should be included in the model. Be that as it may, since (7) holds, the proposed method is theoretically guaranteed to be consistent. Under this setting, we compare the proposed method with a naive method estimating the propensity score by MLE, and confirm whether the proposed method is robust to the misspecification of the propensity score model by numerical experiments. The number of repetitions is 3000 in all numerical experiments of this article.
In Table 1, the true values of the ATT, the mean of the estimates and their empirical confidence intervals are given for several $\alpha^{*}$, $\beta^{*}$ and sample sizes $n$. With increasing $\alpha^{*}$, which represents the degree of the misspecification of the propensity score model, the bias of the naive method becomes larger, and the true value is no longer included in the confidence interval. Regarding the proposed method, it is found that the approximation accuracy is considerably high when the weighting matrix in the generalized method of moments is the identity matrix. On the other hand, when the weighting matrix is an optimal matrix, the accuracy is often comparable to that of the identity matrix, but the accuracy is similar to that of the naive method in some cases. Since there are settings in the simulation where the optimal matrix is close to degenerate, this may be attributed to the fact that the estimation of the $\theta$ is not stable. Regardless of whether or not (7) holds, if the propensity score model is estimated correctly, both the proposed method and the naive method can estimate ATT without bias.
Table 1: Comparison of the estimator of the ATT when the propensity score is given by the CBD method and the MLE: True is the true value of the ATT, CBD-id is the CBD method when the weighting matrix in the generalized method of moments is the identity matrix, and CBD-opt is the CBD method when it is the optimal matrix. Est is the average of the estimators obtained for each simulation, and Conf Int is its empirical 95% confidence interval.
CBD-id CBD-opt MLE
$\beta^{*}$ $\alpha^{*}$ $n$ True Est Conf Int Est Conf Int Est Conf Int
200 0.09 0.09 [--0.12, 0.30] 0.17 [--0.20, 0.51] --0.00 [--0.22, 0.21]
1.0 400 0.08 0.09 [--0.07, 0.24] 0.17 [--0.14, 0.45] --0.01 [--0.17, 0.14]
600 0.08 0.08 [--0.03, 0.21] 0.19 [--0.13, 0.12] --0.01 [--0.13, 0.12]
200 0.10 0.09 [--0.13, 0.32] 0.13 [--0.10, 0.37] --0.14 [--0.43, 0.15]
0.1 2.0 400 0.10 0.10 [--0.08, 0.25] 0.13 [--0.04, 0.29] --0.13 [--0.34, 0.06]
600 0.09 0.10 [--0.04, 0.24] 0.13 [--0.01, 0.27] --0.12 [--0.30, 0.03]
200 0.09 0.11 [--0.15, 0.39] 0.12 [--0.14, 0.37] --0.31 [--0.80, 0.10]
3.0 400 0.09 0.11 [--0.07, 0.29] 0.13 [--0.06, 0.31] --0.30 [--0.60, --0.03]
600 0.10 0.11 [--0.03, 0.26] 0.12 [--0.02, 0.27] --0.30 [--0.55, --0.08]
200 0.42 0.43 [0.21, 0.65] 0.52 [0.09, 0.93] 0.31 [0.06, 0.57]
1.0 400 0.43 0.43 [0.27, 0.59] 0.53 [0.17, 0.85] 0.31 [0.15, 0.49]
600 0.43 0.43 [0.31, 0.56] 0.55 [0.20, 0.80] 0.31 [0.18, 0.45]
200 0.47 0.46 [0.23, 0.70] 0.50 [0.27, 0.75] 0.21 [--0.10, 0.51]
0.5 2.0 400 0.46 0.47 [0.29, 0.63] 0.50 [0.32, 0.67] 0.21 [--0.00, 0.41]
600 0.46 0.47 [0.33, 0.61] 0.51 [0.37, 0.64] 0.22 [0.04, 0.39]
200 0.47 0.50 [0.24, 0.78] 0.51 [0.25, 0.77] 0.05 [--0.45, 0.47]
3.0 400 0.47 0.50 [0.32, 0.69] 0.52 [0.33, 0.70] 0.06 [0.32, 0.69]
600 0.48 0.50 [0.35, 0.66] 0.51 [0.37, 0.66] 0.05 [--0.20, 0.28]
200 0.85 0.85 [0.62, 1.10] 0.97 [0.47, 1.44] 0.71 [0.44, 1.01]
1.0 400 0.85 0.85 [0.68, 1.03] 0.98 [0.53, 1.35] 0.70 [0.52, 0.91]
600 0.85 0.85 [0.72, 0.99] 1.00 [0.57, 1.31] 0.71 [0.55, 0.87]
200 0.92 0.93 [0.68, 1.19] 0.97 [0.73, 1.24] 0.63 [0.30, 0.96]
1.0 2.0 400 0.92 0.93 [0.74, 1.10] 0.97 [0.78, 1.15] 0.64 [0.40, 0.85]
600 0.92 0.93 [0.78, 1.08] 0.98 [0.83, 1.12] 0.65 [0.46, 0.83]
200 0.95 0.99 [0.71, 1.28] 1.00 [0.73, 1.27] 0.50 [--0.02, 0.93]
3.0 400 0.95 0.99 [0.80, 1.18] 1.00 [0.81, 1.20] 0.50 [0.18, 0.80]
600 0.95 0.99 [0.83, 1.15] 1.00 [0.85, 1.16] 0.50 [0.24, 0.74]
200 2.56 2.55 [2.13, 2.96] 2.77 [1.84, 3.49] 2.30 [1.82, 2.82]
1.0 400 2.56 2.55 [2.27, 2.86] 2.78 [1.96, 2.67] 2.30 [1.95, 2.67]
600 2.56 2.55 [2.33, 2.76] 2.82 [2.03, 3.32] 2.30 [2.02, 2.58]
200 2.75 2.78 [2.42, 3.18] 2.86 [2.51, 3.24] 2.35 [1.91, 2.86]
3.0 2.0 400 2.75 2.78 [2.51, 3.04] 2.85 [2.60, 3.11] 2.35 [2.01, 2.68]
600 2.75 2.78 [2.57, 3.00] 2.85 [2.66, 3.06] 2.35 [2.07, 2.63]
200 2.85 2.93 [2.55, 3.33] 2.95 [2.59, 3.32] 2.28 [1.70, 2.85]
3.0 400 2.85 2.93 [2.67, 3.21] 2.95 [2.70, 3.21] 2.28 [1.89, 2.67]
600 2.85 2.93 [2.71, 3.15] 2.95 [2.75, 3.16] 2.28 [1.97, 2.59]
4 Model selection criteria for the semiparametric difference-in-difference method when propensity score is known
4.1 Risk function and its asymptotic evaluation in the semiparametric difference-in-difference method
As mentioned in the Introduction, there are no reasonable model selection criteria even for the basic setting as treated in Abadie (2005), we first derive model selection criteria when the propensity score is known, and then develop model selection criteria when using the proposed CBD method. Specifically, using the ideas of Baba et al. (2017) and Baba and Ninomiya (2021), we derive model selection criteria as asymptotically unbiased estimators of the natural risk function corresponding to the loss function used to obtain SDID estimators. By using these criteria, the dimension of $\bm{\theta}$ is determined and conponent of $\bm{x}$ is selected. As its natural risk function, based on (4), we define a weighted risk function,
$\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{{\rm E}[y^{[1]}_{i}(1)-y^{[0]}_{i}(1)\mid\bm{x}_{i},d^{[1]}_{i}=1]-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}]$
$=\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}]$
$\ \phantom{=}-\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-{\rm E}[y^{[1]}_{i}(1)-y^{[0]}_{i}(1)\mid\bm{x}_{i},d^{[1]}_{i}=1]\}^{2}]$
$\ \phantom{=}+2\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-{\rm E}[y^{[1]}_{i}(1)-y^{[0]}_{i}(1)\mid\bm{x}_{i},d^{[1]}_{i}=1]\}$
[…]
According to the derivation of the conventional $C_{p}$ criterion, we remove the expectation in the first term, ignore the second term which is independent of the models, and asymptotically evaluate the third term. For simplicity, we refer to the third term as the bias when the risk is simply estimated. If we define $\varepsilon_{i}\equiv\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-a(\bm{x}_{i})$ using $a(\bm{x}_{i})$ in (1), then the third term can be written as
$2\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-a(\bm{x}_{i})\}\{\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}-a(\bm{x}_{i})\}]$
$=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}-a(\bm{x}_{i})\}]$
$=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\bm{x}_{i}^{{\rm T}}(\hat{\bm{\theta}}-\bm{\theta}^{*})].$ (17)
The second equality follows from ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})a(\bm{x}_{i})]=0$ and ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\allowbreak\bm{x}^{{\rm T}}_{i}\bm{\theta}^{*}]=0$ since ${\rm E}[\varepsilon_{i}\mid\bm{x}_{i}]=0$ holds from (3). Here, we use the same $e^{[1]}(\bm{x}_{i};\bm{\alpha})$ as that used to construct $\hat{\bm{\theta}}$. Then, letting $c^{\rm limit}$ be the weak limit of the third term with expectation removed, we consider ${\rm E}[c^{\rm limit}]$ as the asymptotic evaluation of the third term of (16). Note that we consider a risk function based on the loss function in deriving the SDID estimator, but it is also possible to consider a simple squared error.
4.2 Derivation of Model Selection Criterion
In the case of known propensity scores, we denote $e^{[1]}(\bm{x}_{i};\bm{\alpha}^{*})=e^{[1]}(\bm{x}_{i})$, $\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha}^{*})=\rho(\bm{t}_{i},\bm{x}_{i})$ for simplicity. The main term of the error of the SDID estimator is given by
$\hat{\bm{\theta}}-\bm{\theta}^{*}$ $={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\{\rho(\bm{t}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\{1+{\rm o}_{\rm P}(1)\}$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\{\varepsilon_{i}+a(\bm{x}_{i})-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\{1+{\rm o}_{\rm P}(1)\}.$
Thus, (17) evaluates to
${\rm E}[c^{\rm limit}]=2\sum_{i=1}^{n}{\rm E}\bigg{[}\varepsilon_{i}e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j})\bm{x}_{j}\{\varepsilon_{j}+a(\bm{x}_{j})-\bm{x}_{j}^{{\rm T}}\bm{\theta}^{*}\}\bigg{]}.$
This expectation for the case in which $i\neq j$ is the product of the expectations for $i$ and $j$ from the independence among samples. The expected value for $i$ can be calculated as
${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}]={\rm E}_{\bm{x}_{i}}[{\rm E}[\varepsilon_{i}\mid\bm{x}_{i}]e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}]=\bm{0}.$ (18)
Therefore, we have only to consider the case $i=j$. Since $a(\bm{x}_{j})-\bm{x}_{j}^{{\rm T}}\bm{\theta}^{*}$ contains only $\bm{x}_{j}$ as a random variable, this term disappears from 18, and the following theorem is obtained.
Theorem 2.
When the propensity score is known,
${\rm E}[c^{\rm limit}]=2{\rm E}[\varepsilon e^{[1]}(\bm{x})\bm{x}^{{\rm T}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}e^{[1]}(\bm{x})\bm{x}\varepsilon]=2{\rm tr}\{{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[\sigma^{2}(\bm{x})e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]\}$
is obtained, where
$\sigma^{2}(\bm{x})\equiv{\rm E}[\varepsilon^{2}\mid\bm{x}]$ $={\rm E}[\{\rho(\bm{t},\bm{x})\Delta\}^{2}\mid\bm{x}]-{\rm E}[\rho(\bm{t},\bm{x})\Delta\mid\bm{x}]^{2}.$ (19)
While empirical evaluation of ${\rm E}[c^{\rm limit}]$ may seem easily possible, addressing the term ${\rm E}[{\rm E}[\rho(\bm{t},\bm{x})\Delta\mid\bm{x}]^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]$ must be considered. The reason is that this cannot be empirically evaluated as $1/n\sum_{i=1}^{n}\rho(\bm{t},\bm{x})^{2}\allowbreak\Delta^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}$. In this case, we will substitute $\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}$ for ${\rm E}[\rho(\bm{t}_{i},\bm{x}_{i})\Delta\mid\bm{x}_{i}]$. As a result, when the propensity score is known, we propose
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\{\rho(\bm{t}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}$
$+2{\rm tr}\biggl{(}\biggl{\{}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr{\}}^{-1}\biggl{[}\sum_{i=1}^{n}\Bigl{\{}\rho(\bm{t}_{i},\bm{x}_{i})^{2}\Delta^{2}-(\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}})^{2}\Bigr{\}}e^{[1]}(\bm{x}_{i})^{2}\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr{]}\biggr{)}$ (20)
as the model selection criterion for the SDID method.
4.3 Extension of the QICw
In order to provide a comparator for the proposed model selection criterion, we extend the QICw (Platt et al. 2013) for the setting of this article, which is the model selection criterion for propensity score weighted estimation on missing data. First, for the term of goodness of fit, we construct it using the same ideas as for the QICw. Specifically, the log-likelihood function with no missing data is differentiated, a weighted potential outcome variable is substituted for the one whose expected value is the same as the outcome variable, the differentiation is reverted, and the estimator is substituted for the parameter. The number of parameters is used as the penalty term. In this setting, the log-likelihood function with no missing data should be
$-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}[\{\mbox{$y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$}\}-\bm{x}_{i}^{{\rm T}}\bm{\theta}]^{2},$
and if we take $\rho_{i}\Delta_{i}$ as the weighted potential outcome variable, the outcome variable and the conditional expectation will be the same from (3). Then, the sum of the goodness of fit and penalty terms multiplied by the constant $2\sigma^{2}$, that is,
$\sum_{i=1}^{n}\{\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}+2\sigma^{2}p,$
we call QICw for SDID. In fact, if $\sigma^{2}(\bm{x})$ is constant in our proposed criterion, then it is equivalent to the QICw. The problem is to estimate the variance $\sigma^{2}$ of “$y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$”. This estimated value cannot be obtained from the data using only the present assumptions. Since $y^{[1]}_{i}(1)-y^{[0]}_{i}(1)=\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ holds from $y^{[0]}(0)=y^{[1]}(0)$, and from Assumption 1, the expected value of “$\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$” is the same as the expected value of “$\Delta_{i}^{[1]}$ in $d^{[0]}_{i}=0$ minus $\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$”. Since the variance must be evaluated, we further assume that $\Delta_{i}^{[1]}$ and $\Delta_{i}^{[0]}$ are independent conditioned on $d^{[1]}=1$ and $d^{[0]}=1$, respectively, and calculate that
$\sigma^{2}$ $={\rm Var}[\Delta^{[1]}\mid d^{[1]}=1]+{\rm Var}[\Delta^{[0]}\mid d^{[0]}=1]$
$={\rm E}[\Delta^{2}\mid d^{[1]}=1]-{\rm E}[\Delta\mid d^{[1]}=1]^{2}+{\rm E}[\Delta^{2}\mid d^{[0]}=1]-{\rm E}[\Delta\mid d^{[0]}=1]^{2}$
From the above, we define $n_{1}\equiv\sum_{i=1}^{n}d^{[1]}_{i}$ and $n_{0}\equiv\sum_{i=1}^{n}d^{[0]}_{i}$, and
$\hat{\sigma}^{2}=\frac{1}{n_{1}}\sum_{i=1}^{n}d^{[1]}_{i}\Delta_{i}^{2}-\bigg{(}\frac{1}{n_{1}}\sum_{i=1}^{n}d^{[1]}_{i}\Delta_{i}\bigg{)}^{2}+\frac{1}{n_{0}}\sum_{i=1}^{n}d^{[0]}_{i}\Delta_{i}^{2}-\bigg{(}\frac{1}{n_{0}}\sum_{i=1}^{n}d^{[0]}_{i}\Delta_{i}\bigg{)}^{2}$
is used in the present study.
Remark 3.
Let us compare $\sigma^{2}(\bm{x})$ in the proposed criterion with $\sigma^{2}$ in the QICw. In (19), if all expected values do not actually depend on $\bm{x}$, then
$\sigma^{2}(\bm{x})=$ ${\rm E}\bigg{[}\frac{\Delta^{[1]2}}{e^{[1]}}\ \bigg{|}\ d^{[1]}=1\bigg{]}-{\rm E}[\Delta^{[1]}\mid d^{[1]}=1]^{2}+{\rm E}\bigg{[}\frac{\Delta^{[0]2}}{e^{[0]}}\ \bigg{|}\ d^{[0]}=1\bigg{]}-{\rm E}[\Delta^{[0]}\mid d^{[0]}=1]^{2}$
$+2{\rm E}[\Delta^{[1]}\mid d^{[1]}=1]{\rm E}[\Delta^{[0]}\mid d^{[0]}=1].$
Since $\Delta^{[1]}$ and $\Delta^{[0]}$ include fluctuations over time, it is natural that $\Delta^{[1]}\Delta^{[0]}$ should be non-negative. In other words, basically $\sigma^{2}(\bm{x})>\sigma^{2}$, and the penalty term of the proposed criterion is larger than that of the QICw.
4.4 Numerical experiments
By simulations, we evaluate the performance of the proposed criterion (20) when the propensity score is known. As data generation settings, we consider the Case 1-1,
$x_{1}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1})),$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+1+\beta^{*}x_{1}+\epsilon^{[1]},$
and the Case 1-2,
$x_{1},x_{2}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})),$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+1+\beta^{*}x_{1}+\epsilon^{[1]}.$
In both cases, $\beta^{*}$ represents the degree of contribution of the covariate $x_{1}$ to the causal effect.
In Table 2, for some $\beta^{*}$ and sample size $n$, we examine whether the second term of (20), the penalty term of the proposed criterion, accurately approximates the bias (17) that we originally wanted to evaluate. In all cases, our penalty terms are close to the true values of the bias, while the penalty terms in the QICw significantly underestimate the bias.
Table 2: Bias evaluation when the propensity score is known: True is the true value of the bias evaluated by the Monte Carlo method, and Proposal is the penalty term of the proposed method. The penalty term of the QICw is also given as a reference.
Case 1-1 Case 1-2
$\beta^{*}$ $n$ True Proposal QICw True Proposal QICw
200  37.54  37.29 2.23  39.20  39.43  5.96
0.1 400  38.56  37.71 2.24  39.75  39.74  5.98
600  37.89  37.89 2.27  39.79  39.68  6.00
200  57.34  56.28 2.31  59.77  58.69  6.39
0.5 400  59.36  56.28 2.33  59.36  59.52  6.48
600  58.56  57.35 2.34  60.20  59.37  6.48
200  92.48  91.33 2.56 103.24  98.85  7.91
1.0 400  98.17  91.43 2.58 108.09  99.46  7.99
600  87.03  92.62 2.59  98.33  99.57  7.99
200 359.64 358.56 5.17 439.35 429.91 23.65
3.0 400 345.73 348.96 5.20 441.47 430.88 24.05
600 367.73 365.75 5.22 435.32 430.09 23.87
Next, we examine the performance of model selection using the proposed criteria. Data is generated from
$x_{1},x_{2},\ldots,x_{l}\sim{\rm Uniform}(0,2),\quad\bm{x}=(1,x_{1},x_{2},\ldots,x_{l})^{{\rm T}},$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+\bm{\beta}^{[1]*}{}^{{\rm T}}\bm{x}+\epsilon^{[1]},$
and we consider the Case 2-1,
$l=4,\quad\bm{\beta}^{[1]*}=(1,\beta^{*},\bm{0}_{3}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1})),$
the Case 2-2,
$l=4,\quad\bm{\beta}^{[1]*}=(1,\bm{1}_{2}^{{\rm T}}\beta^{*},\bm{0}_{2}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})),$
and the Case 2-3,
$l=6,\quad\bm{\beta}^{[1]*}=(1,\bm{1}_{2}^{{\rm T}}\beta^{*},\bm{0}_{4}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})).$
Table 3 compares the proposed criterion and the QICw by selecting models using the forward selection method. As the main index to measure the goodness of the criteria, we use the empirical estimator of the risk function, specifically, the average of 3000 calculations of
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha})(\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}})^{2}.$
In the Case 2-2 setting, where there are relatively few unnecessary covariates among the candidates, that is, the negative effect of false positives is less likely to appear, the QICw slightly exceeds the proposed criterion in some cases for $\beta^{*}=0.5$ and $\beta^{*}=1.0$. Otherwise, the proposed criterion outperforms the QICw, especially in settings with many covariates unrelated to the true structure such as Case 2-1 and Case 2-3, a clear difference is observed. Since the QICw underestimates the penalty term, the false positive value must be large and too many extraneous variables must be selected in all settings.
Table 3: Comparison of the proposed criterion and the QICw through model selection when the propensity score is known: Risk is the empirical estimate of the risk of the selected model, TP is the average number of covariates that are true positives, and FP is the average number of covariates that are false positives.
Case 2-1 Case 2-2 Case 2-3
$\beta^{*}$ $n$ Risk TP FP Risk TP FP Risk TP FP
200 Proposal 11.05 0.50 1.29  14.07 0.61 0.62  19.01 0.63 1.15
QICw 11.29 0.53 1.47  15.94 0.90 0.95  21.74 0.90 1.73
0.1 400 Proposal 11.04 0.47 1.30  13.99 0.68 0.52  19.72 0.69 1.08
QICw 11.27 0.51 1.49  15.23 0.98 0.86  22.45 0.98 1.77
600 Proposal 10.86 0.48 1.25  14.40 0.75 0.51  19.51 0.72 1.14
QICw 11.00 0.51 1.44  15.88 1.07 0.83  22.62 1.02 1.81
200 Proposal 14.68 0.70 1.35  23.20 1.23 0.61  30.01 1.20 1.21
QICw 14.94 0.74 1.61  23.33 1.50 1.00  31.61 1.48 1.96
0.5 400 Proposal 14.42 0.83 1.29  22.95 1.60 0.60  31.82 1.55 1.14
QICw 14.71 0.86 1.58  22.48 1.78 1.01  32.67 1.75 1.93
600 Proposal 14.67 0.89 1.31  23.24 1.74 0.57  30.25 1.75 1.20
QICw 14.93 0.91 1.61  23.00 1.88 0.98  31.96 1.88 1.98
200 Proposal 21.69 0.89 1.39  37.19 1.69 0.61  48.24 1.69 1.26
QICw 22.00 0.91 1.67  36.59 1.85 1.04  49.53 1.84 2.18
1.0 400 Proposal 21.09 0.97 1.40  34.65 1.95 0.63  46.07 1.94 1.20
QICw 21.61 0.98 1.76  35.71 1.99 1.08  49.45 1.98 2.12
600 Proposal 20.74 1.00 1.35  32.16 1.99 0.60  43.66 2.00 1.24
QICw 21.30 1.00 1.72  34.38 1.99 1.07  48.02 2.00 2.19
200 Proposal 67.79 0.99 1.54 136.61 1.96 0.67 180.41 1.96 1.33
QICw 69.22 1.00 1.93 141.09 1.99 1.19 194.02 1.98 2.34
3.0 400 Proposal 66.20 1.00 1.57 133.32 2.00 1.20 175.10 2.00 1.25
QICw 67.83 1.00 1.97 141.74 2.00 1.20 191.79 2.00 2.38
600 Proposal 65.76 1.00 1.48 126.36 2.00 0.62 173.78 2.00 1.29
QICw 67.51 1.00 1.91 135.24 2.00 1.17 191.42 2.00 2.36
5 Model Selection Criteria for the CBD Method
5.1 Derivation of Model Selection Criterion
As in the case with known propensity scores in Section 4.2, we derive the model selection criteria when using the CBD method. We evaluate
$\sum_{i=1}^{n}2{\rm E}[e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{t}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-{\rm E}[y^{[1]}_{i}-y^{[0]}_{i}\mid\bm{x}_{i},d^{[1]}_{i}=1]\}\bm{x}_{i}^{{\rm T}}(\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*})],$
which is the third term of the risk function (17). First, substituting $\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}$ for ${\rm E}[y^{[1]}_{i}-y^{[0]}_{i}\mid\bm{x}_{i},d^{[1]}_{i}=1]$, and then substituting the right-hand side of (15) for $\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$, this third term can be written as
$2{\rm E}\bigg{[}\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{t}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}^{{\rm T}}$
[…]
Since $(1/n)\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{j}\bm{x}_{j}^{{\rm T}}$ converges to
$\bm{L}(\bm{\alpha}^{\dagger})\equiv{\rm E}\{e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\bm{x}\bm{x}^{{\rm T}}\}$
in probability, if we let $\bm{u}$ be the normal random vector that is the weak limit of
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i},$ (21)
then the weak limit of (17) removing the expectation is expressed as
$c^{\rm limit}=2\bm{u}^{{\rm T}}\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{u}=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{u}\bm{u}^{{\rm T}}\}.$
In the setting where the estimator $\hat{\bm{\theta}}^{\rm CBD}$ is consistent with $\bm{\theta}^{*}$, ${\rm E}[\bm{u}\bm{u}^{{\rm T}}]={\rm Var}[\bm{u}]$ holds because the expected value of $\bm{u}$ becomes $\bm{0}$, and the asymptotic evaluation of (17), which is to be used in the information criterion, is expressed as
${\rm E}[c^{\rm limit}]=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}{\rm Var}[\bm{u}]\}.$
Here, we write down ${\rm Var}[\bm{u}]$. As the error of the parameter $\hat{\bm{\alpha}}^{\rm CB}$ in the propensity score can be shown to be (14), using this, (21) can be expanded as
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\bm{x}_{i}\bigg{\{}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}$
$\ \ \ \ \ +\dfrac{\partial}{\partial\bm{\alpha}^{{\rm T}}}\bigg{(}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bigg{[}\frac{d^{[1]}_{i}-e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{1-e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\}}\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\bigg{]}\bigg{)}(\hat{\bm{\alpha}}^{\rm CB}-\bm{\alpha}^{\dagger})\bigg{\}}\{1+{\rm o}_{\rm P}(1)\}$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\Big{[}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}$
$\ \ \ \ \ \ \ \ \ \ -\bm{M}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\}^{-1}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Big{]}\{1+{\rm o}_{\rm P}(1)\},$
where
$\bm{M}(\bm{\alpha},\bm{\theta})={\rm E}\biggl{[}\dfrac{\partial}{\partial\bm{\alpha}}e^{[1]}(\bm{x};\bm{\alpha})\biggl{[}\frac{d^{[1]}-1}{\{1-e^{[1]}(\bm{x};\bm{\alpha})\}^{2}}\Delta-\bm{x}^{{\rm T}}\bm{\theta}\biggr{]}\bm{x}\biggr{]}.$
From the above, by defining
$\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\equiv{\rm Var}$ $[e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\{\rho(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}$
$-\bm{M}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\}^{-1}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})],$ (22)
it can be shown to be ${\rm Var}[\bm{u}]=\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$. Thus, the following theorem is obtained.
Theorem 3.
When the propensity score is estimated by the CBD method, the third term of (16) is evaluated as
${\rm E}[c^{\rm limit}]=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\}.$
As a result, when the propensity score is estimated by the CBD method, we propose
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})(\rho(\bm{d}_{i},\bm{x}_{i})(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD})^{2}+2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\}$ (23)
as the model selection criterion for the SDID method, which is an asymptotically unbiased estimator of (16).
Also in this case, the penalty term depends on the true values of the parameters, thus empirical estimates are used in practice. Specifically, $\bm{L}_{n}(\hat{\bm{\alpha}}^{\rm CB})\equiv n^{-1}\sum_{i=1}^{n}\allowbreak e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}$ is used instead of $\bm{L}(\bm{\alpha}^{\dagger})$ and
$\bm{V}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})\equiv\frac{1}{n}$ $\sum_{i=1}^{n}\Bigl{[}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD}\}\bm{x}_{i}$
$-\bm{M}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})\{\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{G}_{n}(\hat{\bm{\alpha}}^{\rm CB})\}^{-1}\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Bigr{]}^{2}$
is used instead of $\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$, where
$\bm{M}_{n}(\bm{\alpha},\bm{\theta})\equiv\frac{1}{n}\sum_{i=1}^{n}\dfrac{\partial}{\partial\bm{\alpha}}e^{[1]}(\bm{x}_{i};\bm{\alpha})\biggl{(}\frac{d^{[1]}_{i}-1}{(1-e^{[1]}(\bm{x}_{i};\bm{\alpha}))^{2}}\Delta_{i}-\bm{x_{i}}^{{\rm T}}\bm{\theta}\biggr{)}\bm{x_{i}}.$
Here, the propensity score is estimated using the covariate balancing approach, but the information criterion can be derived in the same way when the MLE is used (see Supplementary material).
5.2 Numerical experiments
Through simulations, we examine the accuracy of the approximation of the penalty term and the performance of the model selection criterion (23) based on Theorem 3 for the proposed CBD method. As the weighting matrix in the generalized method of moments, we use the identity matrix based on the results of Section 4.4. But even when an optimal matrix is used as the weighting matrix, the accuracy of the approximation of the penalty term and the performance of the model selection are comparable to those obtained when the identity matrix is used (see Supplementary Material).
In Table 4, data are generated as in Case 1-1 and Case 1-2 of Section 4.4 to verify whether the second term of (23), the penalty term of the proposed criterion, accurately approximates the bias (17) that we originally wanted to evaluate for some parameter values $\beta^{*}$ and sample size $n$. In all cases, the penalty term of the proposed criterion provides values much closer to the bias than the penalty term of the QICw, and its approximation accuracy is confirmed to be high. Conversely, in all cases, the penalty term of the QICw significantly underestimates the bias.
Table 4: Bias evaluation when the propensity score is estimated by the CBD method with identity matrix as the weighting matrix: True is the true value of the bias evaluated by the Monte Carlo method, and Proposal is the penalty term of the proposed method. The penalty term of the QICw is also given as a reference.
Case 1-1 Case 1-2
$\beta^{*}$ $n$ True Proposal QICw True Proposal QICw
200  7.53  7.35 2.23 17.61 16.98  5.94
0.1 400  7.19  7.31 2.24 17.88 17.55  5.99
600  7.09  7.34 2.27 17.10 17.35  5.98
200  8.21  8.04 2.31 19.42 18.76  6.38
0.5 400  7.79  7.89 2.33 19.51 19.62  6.48
600  7.75  7.88 2.34 18.94 19.25  6.47
200  9.47  9.56 2.58 23.20 22.67  7.92
1.0 400  9.04  9.25 2.58 23.07 23.18  8.00
600  8.96  9.14 2.59 22.62 22.98  8.00
200 19.10 19.22 5.17 54.55 54.82 23.62
3.0 400 18.36 18.58 5.20 53.71 54.81 24.08
600 17.88 18.49 5.22 53.55 54.64 23.90
In Table 5, data are generated as in Case 2-1, Case 2-2 and Case 2-3 of Section 4.4 and the proposed criterion and the QICw are compared by model selection using the forward selection method. The main index of the goodness of the model selection criteria is again the empirical estimate of the risk. Unlike the results in Section 4.4, where the propensity score is known, the proposed criterion always outperforms the QICw, even in Case 2-2, where relatively few covariates are unnecessary among the candidates. On the other hand, as in Case 2-1 and Case 2-3, the proposed criterion is particularly superior when the candidates include many covariates that do not contribute to the true structure, which is similar to Section 4.4. The penalty term is underestimated in the QICw, therefore, too many extraneous variables must be selected.
Table 5: Comparison of the proposed criterion and the QICw through model selection when the propensity score is estimated using the CBD method with identity matrix as the weighting matrix: Risk is the empirical estimate of the risk of the selected model, TP is the average number of covariates that are true positives, and FP is the average number of covariates that are false positives.
Case 2-1 Case 2-2 Case 2-3
$\beta^{*}$ $n$ Risk TP FP Risk TP FP Risk TP FP
200 Proposal 7.33 0.21 0.53 12.38 0.41 0.33 16.49 0.38 0.74
QICw 10.32 0.48 1.55 16.96 0.91 0.92 22.57 0.84 1.79
0.1 400 Proposal 7.30 0.21 0.49 12.06 0.43 0.33 17.02 0.46 0.69
QICw 10.13 0.47 1.46 15.75 0.96 0.91 22.47 0.96 1.74
600 Proposal 6.88 0.20 0.46 12.01 0.46 0.34 16.62 0.50 0.66
QICw 9.57 0.50 1.43 15.11 0.96 0.91 21.91 0.98 1.79
200 Proposal 10.73 0.51 0.55 20.12 1.12 0.34 26.11 1.11 0.72
QICw 13.12 0.77 1.65 21.56 1.55 1.01 30.24 1.51 2.01
0.5 400 Proposal 10.64 0.72 0.50 20.34 1.52 0.33 26.36 1.54 0.69
QICw 12.87 0.88 1.61 20.47 1.80 1.02 29.42 1.80 1.97
600 Proposal 10.02 0.84 0.48 17.98 1.79 0.33 24.29 1.77 0.66.
QICw 12.05 0.97 1.60 19.15 1.92 1.03 28.46 1.92 2.01
200 Proposal 13.83 0.86 0.54 25.04 1.78 0.33 34.29 1.80 0.72
QICw 17.55 0.95 1.78 28.00 1.92 1.10 41.46 1.93 2.15
1.0 400 Proposal 12.44 0.98 0.50 22.03 1.97 0.33 31.09 1.97 0.67
QICw 16.82 1.00 1.76 26.42 2.00 1.10 40.19 2.00 2.16
600 Proposal 11.33 1.00 0.48 20.06 2.00 0.31 28.79 2.00 0.63
QICw 15.91 1.00 1.76 25.28 2.00 1.11 39.30 2.00 2.17
200 Proposal 34.16 1.00 0.52 60.56 2.00 0.31 95.78 2.00 0.68
QICw 48.77 1.00 1.94 81.22 2.00 1.17 136.38 2.00 2.37
3.0 400 Proposal 31.96 1.00 0.51 59.75 2.00 0.31 91.38 2.00 0.65
QICw 46.25 1.00 1.95 80.71 2.00 1.18 132.43 2.00 2.36
600 Proposal 29.56 1.00 0.49 57.25 2.00 0.31 87.98 2.00 0.61
QICw 44.02 1.00 1.95 78.45 2.00 1.18 130.44 2.00 2.34
6 Real data analysis
The LaLonde dataset is treated in LaLonde (1986) and included in the R package Matching version 4.10.14 (Sekhon 2011). The group that took the U.S. job training program in 1976 is denoted as $t=1$ and the group that did not take the program is denoted as $t=0$. The difference in annual income in 1978 after the training for each group is estimated via the ATT. Age (age, years of education (educ), income in 1974 (re74), black (black), hispanic (hisp), married (married), and high school graduate or higher (nodegr) are treated as covariates, i.e. $p=7$, and the outcome variable is the difference between income in 1978 (re78) and income in 1974 (re78). The sample size is $N=445$. The objective of this section is not to extract information from this data, but to examine whether the proposed method makes a significant difference from existing methods, and therefore, a sufficiently small number of samples are assumed. Specifically, for a natural number $k$, the $23k-$ 2 rows of data are analyzed in Block 1, $13k-$ 1 rows in Block 2, and $3k$ rows in Block 3, respectively.
Regarding the estimation, we first estimate the propensity score using the CBD method with the identity matrix as the weighting matrix. Then, assuming that the ATT can be written as a linear sum of covariates, we estimate its coefficients by the SDID method. Table 5 selects covariates using the proposed criterion or the QICw with the forward selection method and estimates the coefficients of the selected variables. The QICw selects all variables in any Block, while the proposed criterion does not select some variables. In particular, the difference between the models selected for Block 1 and Block 3 is remarkable. Because we do not know the true structure of the real data, it is impossible to judge the superiority or inferiority of the two criteria, however, we can see that the difference between them is quite large. In a setting such as this article, it is essential to use model selection criteria that are theoretically valid rather than intuitive ones.
Table 6: Comparison of the proposed criterion and the QICw when propensity scores are estimated using the CBD method with the identity matrix as the weighing matrix, through model selection experiments on LaLonde data. The numbers with exponential notation are the estimated coefficients for each variable, and a 00 means that the variable was not selected in the model selection.
Block intercept age educ re74
1 Proposal --2.14E+3 0E+0 0E+0 0E+0
QICw --1.37E+4 5.29E+2 4.43E+1 --2.01E-1
2 Proposal 1.59E+4 0E+0 1.08E+3 2.54E-1
QICw 2.38E+4 --1.91E+2 8.22E+2 2.73E-1
3 Proposal --8.50E+2 0E+0 0E+0 0E+0
QICw --8.51E+3 6.13E+0 3.35E+2 2.56E-1
Block black hisp married nodegr
1 Proposal 0E+0 0E+0 0E+0 0E+0
QICw 3.55E+2 --8.19E+3 --9.41E+2 --1.30E+3
2 Proposal --1.86E+4 --2.34E+4 4.02E+3 --1.26E+4
QICw --1.85E+4 --2.42E+4 4.88E+3 --1.34E+4
3 Proposal 0E+0 --1.03E+4 0E+0 0E+0
QICw 3.84E+3 --6.71E+3 2.92E+3 --5.90E+2
7 Discussion
In this article, we have proposed the CBD method, a SDID method incorporating covariate balancing. In order to ensure the doubly robustness, i.e., that the ATT can be consistently estimated if the model for the propensity score is correctly specified or the model for the changes in the outcomes over time is correctly assumed, we have shown that it is sufficient to balance the distribution up to the second-order moment of the covariates across groups with the treatment group as the target population. This fact is important to note because covariate balancing usually balances first-order moment. We also have developed model selection criteria for the CBD method as an asymptotically unbiased estimator of the risk based on the loss function used for the estimation. Specifically, we have derived model selection criteria using a DID-specific assumption called the conditional parallel trend assumption, i.e., without relying on the ignorable treatment assignment condition, which is often assumed in causal inference, and we have shown that the penalty term is considerably different from twice the number of parameters that often appears in AIC-type model selection criteria. Since there have been no model selection criteria for even the basic SDID method, we have developed the model selection criteria not only for the CBD method, but also for when the propensity score is known or the propensity score is estimated by the MLE. Numerical experiments showed that the proposed methods estimate ATT robustly compared to the method that estimates propensity scores by the MLE, and that the proposed criteria clearly reduce the risk targeted by the SDID method compared to the intuitive generalization of the existing information criterion to the SDID method. In addition, the real data analysis has confirmed that there were large differences between the results of the proposed method and the existing methods.
Although this article deals with a linear model of covariates as a simple working model for the outcome, it should be noted that the results can be easily extended to situations where a nonlinear model of covariates should be used, for example, when using the kernel method. In fact, Hazlett (2020) treats the function $\phi(\bm{x}):\mathbb{R}^{p}\rightarrow\mathbb{R}^{q}$ in the reproducing kernel Hilbert space as $q$ may be larger than $p$ or $n$. And they assume that the outcome is represented by a linear model of $\phi(\bm{x})$, the feature map of $\bm{x}$, which in our setting can be written as
${\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]=\phi(\bm{x})^{{\rm T}}\bm{\theta}.$ (24)
Then, by using
$\sum_{i=1}^{n}\bigg{\{}\frac{d^{[1]}_{i}\phi(\bm{x}_{i})}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-\frac{d^{[0]}_{i}\phi(\bm{x}_{i})}{e^{[0]}(\bm{x}_{i};\bm{\alpha})}\bigg{\}}=0$
as the moment condition for estimating the propensity score, they propose the kernel balancing method, which can flexibly estimate causal effects even if the relationship between the covariate and the outcome is nonlinear. In the setting of this article, we also consider the situation (24) and when making the estimation doubly robust, we assume that the change in the outcomes follows
${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\phi(\bm{x})^{{\rm T}}\bm{\beta}^{[k]*}.$
Let $k(\cdot,\cdot):\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$ be the kernel function corresponding to $\phi(\bm{x})$, and let $\bm{K}$ be a $p\times p$ semidefinite symmetric matrix whose $(l,m)$ component is $k(x_{l},x_{m})$. If we estimate the parameter $\bm{\alpha}$ that satisfies
$\bm{H}^{[1]}(\bm{d},\bm{x};\bm{\alpha})={\rm E}\bigg{[}e^{[1]}(\bm{x};\bm{\alpha})\bigg{\{}\frac{d^{[1]}}{e^{[1]}(\bm{x};\bm{\alpha})}-1\bigg{\}}\bm{K}\bigg{]}=\bm{{\rm O}},$
$\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})={\rm E}\bigg{[}e^{[1]}(\bm{x};\bm{\alpha})\bigg{\{}\frac{d^{[0]}}{e^{[0]}(\bm{x};\bm{\alpha})}-1\bigg{\}}\bm{K}\bigg{]}=\bm{{\rm O}},$
and then estimate $\bm{\theta}$ or derive model selection criteria based on it, that would be a natural extension of the proposed method in this article.
There are other valuable directions for extension. For example, Callaway and Sant’Anna (2021) and Goodman-Bacon (2021) propose a DID method in a more flexible setting than the previous designs, where there are multiple time points and the treatment may be initiated at any of them. Moreover, De Chaisemartin and D’haultfœuille (2023) develops a DID method when there are more than two groups, and Richardson et al. (2023) proposes the generalized DID to estimate causal effects under alternative assumptions to the parallel trend assumption. While the outcome in the basic DID method is continuous, Tchetgen et al. (2024) deals with the case where the response follows a generalized linear model. Thus, the DID method has been theoretically developed in accordance with real problems, and extensions of the proposed method to these problems are recent challenges. In econometrics, the synthetic control method (Abadie et al. 2010) is a major method for causal inference as well as the DID method, and a method combining it with the DID method has also been developed as synthetic DID (Arkhangelsky et al. 2021). In this method, the objective of the information criterion is what to synthesize and with what weights, and the proposed method cannot necessarily be extended in a similar way, but it is an important topic. Furthermore, propensity score analysis with high-dimensional confounding variables is also a hot topic (e.g., Belloni et al. 2017, Chernozhukov et al. 2018, Ning et al. 2020b), and even in that analysis, robustly estimating the influence of covariates that directly affect the causal effect and selecting covariates that truly affect the causal effect is an essential task to ensure the validity of the analysis. The asymptotic properties of the estimator of causal effects tend to be obtained even when the covariates are high-dimensional as well, and it is surely possible to extend our proposed method for these analyses.