### v3

# Covariate balancing estimation and model selection for difference-in-differences approach

Abstract
Remarkable progress has been made in difference-in-differences (DID) approaches to causal inference that estimate the average effect of a treatment on the treated (ATT). Of these, the semiparametric DID (SDID) approach incorporates a propensity score analysis into the DID setup. Supposing that the ATT is a function of covariates, we estimate it by weighting the inverse of the propensity score. In this study, as one way to make the estimation robust to the propensity score modeling, we incorporate covariate balancing. Then, by attentively constructing the moment conditions used in the covariate balancing, we show that the proposed estimator is doubly robust. In addition to the estimation, we also address model selection. In practice, covariate selection is an essential task in statistical analysis, but even in the basic setting of the SDID approach, there are no reasonable information criteria. Here, we derive a model selection criterion as an asymptotically bias-corrected estimator of risk based on the loss function used in the SDID estimation. We show that a penalty term can be derived that is considerably different from almost twice the number of parameters that often appears in AIC-type information criteria.
Keywords: Average effect of treatment on the treated, Doubly robustness, Information criterion, Propensity score analysis, Statistical asymptotic theory
1 Introduction
The difference-in-differences (DID) approach is one of the most common methods used in fields such as economics and epidemiology to assess the effect of a particular intervention or treatment. In the basic setting, there are two groups: a treatment group receiving a specific treatment and a control group receiving no treatment or other treatment. The pre- and post-treatment outcomes of each group are measured and the change in the outcome for the control group is subtracted from the change in the outcome for the treatment group to estimate the average treatment effect on the treated (ATT). Under the parallel trend assumption, an estimation in the DID approach is valid. Recently, Abadie (2005) dealt with inference under a more natural assumption that involves using information from the covariates, specifically a parallel trend assumption conditioned on the covariates. They further proposed a semiparametric DID (SDID) approach that does not necessarily require modeling of the covariates to the outcome and that estimates the ATT by weighting the inverse of the propensity score.
With or without conditionality, the parallel trend assumption is key to the DID approach. This assumption is one where the mean changes over time in the potential outcomes of the treatment group and the control group would have been similar if there had been no treatment. On the other hand, especially in epidemiology, the ignorable treatment assignment assumption has usually been invoked in causal inference (e.g. Robins et al. 1994). This is the assumption that the potential outcome variable and the treatment assignment variable, which represents the received treatment, become independent conditioned on the covariates. Although neither assumption is verifiable, to evaluate the validity of the parallel trend assumption, visual inspection is useful when observations over many time points are available. Moreover, the parallel trend assumption naturally holds if it can be supposed that there is no change over time in the expectation of outcomes without treatment. Needless to say, the parallel trend assumption is considerably weaker than the ignorable treatment assignment assumption, and estimating the ATT under such a desirable assumption is the advantage that the DID approach gains in exchange for collecting pre- and post-treatment outcomes.
The methodological development of the DID approach has been growing because of the above-mentioned advantages and the increasing amount of available observational study data (see Roth et al. 2023 for a comprehensive review of recent studies on the DID approach). However, for example, Abadie (2005) assumes that the model for the propensity score is correctly specified. However, the estimator will have a bias if the model is misspecified. As an extension of Abadie (2005), Sant’Anna and Zhao (2020) assumes not only a regression model for the treatment assignment variable used to obtain the propensity score, but also a regression model for the outcome variable, and proposes an estimator for the ATT derived from the two models. Their estimator is consistent if either the treatment assignment model or the outcome regression model is correctly specified, that is, the estimator is doubly robust.
In this article, as with Abadie (2005), we suppose that the ATT is a function of covariates and address the semiparametric estimation problem for that function. From the estimated function, the heterogeneity of the ATT, that is, the conditional ATT can be evaluated. Two regression models as above are considered, and in order to give double robustness to the estimation, we incorporate covariate balancing (e.g. Egel et al. 2008, Imai and Ratkovic 2014 and Fan et al. 2022), a method using propensity scores such that a certain moment condition is satisfied. This method actually estimates the treatment assignment model, but only supposes an outcome model and does not estimate it. Nevertheless, the SDID estimator derived by weighting the inverse of the propensity score, which is obtained by the covariate balancing, will be consistent if either of the regression models is correctly specified. One point to note is that, in order for the SDID approach to have this favorable asymptotic property, the moment condition has to be constructed in a different way from the usual one, which is due to considering the conditional ATT. Also, the incorporation of covariate balancing leads to an estimation via weights that always balances the moments of the covariates across groups, which is also preferable from an interpretive point of view (Imai and Ratkovic 2014). We call the proposed method covariate balancing for DID (CBD).
Obviously, model selection is also an essential task in using the DID approach in a series of data-analysis tasks. Specifically, by selecting the function of covariates for the ATT addressed in this paper, it becomes possible to evaluate the heterogeneity of the ATT. There have been various important studies conducted in this area. For example, in a study estimating the effect of union membership on wages, Houngbedji (2016) treated the workers’ demographics with a linear model for causal effects and explored factors affecting the causal effect on the basis of the values of the regression coefficients and its $p$-value. In other words, they addressed the issue of covariate selection. However, there is no satisfactory information criterion, a basic tool for model selection, even for the basic setting of SDID. To remedy this situation, we will derive an information criterion for a basic SDID approach and the CBD method that is an asymptotically unbiased estimator of the risk based on the loss function used in the estimation. Note that traditional information criteria such as the generalized information criterion (GIC) (Konishi and Kitagawa 1996) cannot be applied because the loss function includes weights based on the propensity score, which are random variables. Since the SDID approach places importance on being based on the conditional parallel trend assumption, it does not rely on the ignorable treatment assignment assumption when deriving the information criterion, which causes some difficulties. As a result, we will derive a penalty term that is considerably different from twice the number of parameters, which often appears in AIC-type information criteria including the GIC.
In summary, the aim of this paper is to develop a new doubly robust estimation method with the advantage of covariate balancing in the SDID approach and to derive a model selection criterion for the SDID approach, for which no reasonable model selection criteria exist yet. Its organization is as follows. Section 2 describes the model and the estimation target and introduces the SDID approach that was proposed in Abadie (2005). In Section 3, we propose the CBD method and give a theorem showing its double robustness. Numerical experiments indicate that the estimation is robust against model misspecification. In Section 4, we derive a model selection criterion for when the propensity score is known as a bias-corrected estimator of the risk that is naturally defined from the loss function used in the SDID estimation. Then, we also prove a theorem that suggests that the penalty term of the model selection criterion is asymptotically valid. As a comparison, we derive a model selection criterion that naturally extends the idea of QICw in Platt et al. (2013) to the SDID approach. Furthermore, through numerical experiments, we verify the approximation accuracy of the penalty term of the proposed criterion from the perspective of risk estimation and show its superior performance to the QICw from the viewpoint of risk minimization. Section 5 summarizes the results for the CBD method in the same way as Section 4. In other words, a model selection criterion is derived from the naturally defined risk, a theorem is given to support the validity of its penalty term, and its performance is evaluated through numerical experiments. Section 6 reports the results of the estimation and model selection for the real data in LaLonde (1986). Finally, Section 7 discusses future developments of the proposed approach.
2 Preparation
2.1 Model and Assumptions
Suppose that there are two groups, a treatment group and a control group, and that outcomes are collected before and after the possible treatment time point, respectively. We consider the counterfactual setting, where $y^{[1]}(t)$ denotes the outcome variable of the treatment group, $y^{[0]}(t)$ the outcome variable of the control group, $t=0$ before treatment and $t=1$ after treatment. In practice, only one of the outcomes of the treatment group or control group can be obtained. By preparing treatment assignment variables such that $d^{[1]}=1\ (d^{[0]}=0)$ if the sample is assigned to the treatment group and $d^{[0]}=1\ (d^{[1]}=0)$ if the sample is assigned to the control group, the observed outcome variable collected at each time point is expressed as
$y(t)=d^{[1]}y^{[1]}(t)+d^{[0]}y^{[0]}(t).$
Denoting the change in the outcome variables at two time points in the treatment group as $\Delta^{[1]}=y^{[1]}(1)-y^{[1]}(0)$ and the change in the control group as $\Delta^{[0]}=y^{[0]}(1)-y^{[0]}(0)$, the observed change is expressed as
$\Delta=d^{[1]}\Delta^{[1]}+d^{[0]}\Delta^{[0]}.$
Here, for the ATT conditioned on the covariate $\bm{x}\ (\in\mathbb{R}^{p})$,
$a(\bm{x})\equiv{\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1],$ (1)
we suppose $g(\bm{x};\bm{\theta})$ $(\bm{\theta}\in\bm{\Theta}\subset\mathbb{R}^{p})$ as a working model to approximate it, and for simplicity, set the working model to be a linear one $\bm{x}^{{\rm T}}\bm{\theta}$. Then, we estimate $\bm{\theta}$ and give the ATT by taking the expectation with respect to $\bm{x}$ in the treatment group.
Let us define the optimal value of $\bm{\theta}$ as
$\bm{\theta}^{*}\equiv\mathop{\rm argmin}_{\bm{\theta}}{\rm E}[\{{\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]-\bm{x}^{{\rm T}}\bm{\theta}\}^{2}\mid d^{[1]}=1].$ (2)
Abadie (2005) shows that $\bm{\theta}^{*}$ can be consistently estimated with the SDID approach by using propensity scores $e^{[1]}(\bm{x})\equiv{\rm P}(d^{[1]}=1\mid\bm{x})$ and $e^{[0]}(\bm{x})\equiv{\rm P}(d^{[0]}=1\mid\bm{x})$ under several assumptions, and therefore the ATT can be consistently estimated. The assumptions are as follows.
Assumption 1 (Conditional parallel trend).
It holds that
${\rm E}[y^{[0]}(1)-y^{[0]}(0)\mid\bm{x},d^{[1]}=1]={\rm E}[y^{[0]}(1)-y^{[0]}(0)\mid\bm{x},d^{[0]}=1],$
which means that, under the expectation conditioned on $\bm{x}$, the change in outcomes without treatment is the same regardless of whether it is the treatment group or the control group.
Assumption 2 (Common shocks).
Events that occur between the time points of outcome collection before and after treatment have the same impact on outcomes in both the treatment group and the control group.
Assumption 3 (Positivity).
It holds that
$0<{\rm P}(d^{[1]}=1\mid\bm{x})<1;$
that is, the propensity score is not 0 or $1$.
Abadie (2005) showed that, by using weights $\rho(\bm{d},\bm{x})\equiv d^{[1]}/e^{[1]}(\bm{x})-d^{[0]}/e^{[0]}(\bm{x})$,
${\rm E}[\rho(\bm{d},\bm{x})\Delta\mid\bm{x}]={\rm E}[\Delta^{[1]}\mid\bm{x},d^{[1]}=1]-{\rm E}[\Delta^{[0]}\mid\bm{x},d^{[0]}=1]$
$={\rm E}[\Delta^{[1]}-\Delta^{[0]}\mid\bm{x},d^{[1]}=1]={\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]=a(\bm{x})$ (3)
holds. Assumption 1 is used in the second equality. This equation means that, by multiplying the observed values of change with weights based on propensity scores, the conditional ATT can be estimated via the DID approach. Then, from (2) and (3), the optimal solution of $\bm{\theta}$ can be expressed as
$\bm{\theta}^{*}=\mathop{\rm argmin}_{\bm{\theta}}{\rm E}[e^{[1]}(\bm{x})\{\rho(\bm{d},\bm{x})\Delta-\bm{x}^{{\rm T}}\bm{\theta}\}^{2}].$ (4)
Moreover, from this equation,
${\rm E}[e^{[1]}(\bm{x})\bm{x}\{\rho(\bm{d},\bm{x})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\{{\rm E}[\rho(\bm{d},\bm{x})\Delta\mid\bm{x}]-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]={\rm E}[e^{[1]}(\bm{x})\bm{x}\{a(\bm{x})-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]=\bm{0}$ (5)
is satisfied, and therefore,
${\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]\bm{\theta}^{*}$ (6)
is obtained.
2.2 Semiparametric estimation
Here, for an independently and identically distributed set of samples of size $n$, the $i$-th sample will be denoted by the subscript $i$. Abadie (2005) proposed
$\hat{\bm{\theta}}=\bigg\{\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg\}^{-1}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}$ (7)
as a natural estimator of $\bm{\theta}^{*}$ by solving (4). When the observed values of change are multiplied by weights based on propensity scores, a term corresponding to the difference-in-differences appears; thus, this method is called a semiparametric DID (SDID) estimator. The following regularity conditions are assumed.
Assumption 4.
$(\mathrm{i})$ $\bm{\theta}^{*}$ is in the interior of the compact set $\bm{\Theta}\subset\mathbb{R}^{r}$. $(\mathrm{ii})$ ${\rm E}[y(t)^{2}]<\infty$ for all $t$. $(\mathrm{iii})$ ${\rm E}[\bm{x}]$ is bounded and ${\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]$ is non-singular.
From Assumption 4 and (6), we obtain
$\hat{\bm{\theta}}=\bigg\{\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg\}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}$
$\stackrel{{\scriptstyle p}}{{\to}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\rho(\bm{d},\bm{x})\Delta]$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]\bm{\theta}^{*}=\bm{\theta}^{*},$
[…]
${\rm P}(d^{[1]}=1\mid\bm{x})={\rm logit}(\bm{x}^{{\rm T}}\bm{\alpha}^{*}),$
and if $\hat{\bm{\alpha}}$ converges to $\bm{\alpha}^{*}$ under some further assumptions, then $\hat{\bm{\theta}}$ is still a consistent estimator of $\bm{\theta}^{*}$.
3 Covariate balancing for difference-in-differences approach
3.1 Estimation method and its statistical properties
Propensity scores are often estimated by using MLE, but if the model for the propensity scores is misspecified, the estimator of the ATT will have bias. Here, we propose a robust method that can deal with model misspecification. While a doubly robust estimation known as augmented inverse-probability-weighted estimation, which uses estimators from an outcome regression model, has been addressed in Sant’Anna and Zhao (2020) and Ning et al. (2020a), here we construct a doubly robust estimation by incorporating covariate balancing (e.g. Egel et al. 2008, Imai and Ratkovic 2014 and Fan et al. 2022). Let $\bm{{\rm O}}$ be a zero matrix and $\bm{d}=(d^{[0]},d^{[1]})$, and let us consider the following moment conditions:
$\bm{H}^{[1]}(\bm{d},\bm{x};\bm{\alpha})\equiv{\rm E}\bigg[e^{[1]}(\bm{x};\bm{\alpha})\bigg\{\frac{d^{[1]}}{e^{[1]}(\bm{x};\bm{\alpha})}-1\bigg\}\bm{x}\bm{x}^{{\rm T}}\bigg]=\bm{{\rm O}},$
$\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})\equiv{\rm E}\bigg[e^{[1]}(\bm{x};\bm{\alpha})\bigg\{\frac{d^{[0]}}{e^{[0]}(\bm{x};\bm{\alpha})}-1\bigg\}\bm{x}\bm{x}^{{\rm T}}\bigg]=\bm{{\rm O}}.$
From these conditions is obtained ${\rm E}[d^{[1]}\bm{x}\bm{x}^{{\rm T}}]={\rm E}[\{e^{[1]}(\bm{x};\bm{\alpha})/e^{[0]}(\bm{x};\bm{\alpha})\}d^{[0]}\bm{x}\bm{x}^{{\rm T}}]$, which means that the expectation of $\bm{x}\bm{x}^{{\rm T}}$ in the treatment group is adjusted to match the expectation of $\bm{x}\bm{x}^{{\rm T}}$ in the control group through weighting based on the propensity scores. In the original covariate balancing, the first-order moments of the covariates are balanced; that is, $\bm{x}\bm{x}^{{\rm T}}$ in the expectation is naturally set to $\bm{x}$. In this paper, the second-order moments of the covariates are balanced.
The semi-vectorization of the upper triangular part of the matrix is written as ${\rm vech}(\cdot)$. Since $\bm{x}\bm{x}^{{\rm T}}$ is a symmetric matrix, by letting
$\bm{h}(\bm{d},\bm{x};\bm{\alpha})\equiv({\rm vech}\{\bm{H}^{[1]}(\bm{d},\bm{x};\bm{\alpha})\}^{{\rm T}},{\rm vech}\{\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})\}^{{\rm T}})^{{\rm T}},$
the moment conditions can be rewritten as
${\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})]=\bm{0}.$ (8)
We will denote an $\bm{\alpha}$ satisfying (8) as $\bm{\alpha}^{\dagger}$. Since the propensity score model may be misspecified, $\bm{\alpha}^{\dagger}$ is not necessarily the true value $\bm{\alpha}^{*}$, but when the propensity score model is correctly specified, $\bm{\alpha}^{\dagger}=\bm{\alpha}^{*}$. In practice, the SDID approach uses a $\hat{\bm{\alpha}}$ satisfying the empirical version of the moment conditions,
$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})=\bm{0}.$ (9)
If the moment conditions are just-identified, that is, the dimension of the moment conditions (the number of constraints) is the same as the dimension of the parameter $\bm{\alpha}$, then $\bm{\alpha}$ that satisfies (9) exists. On the other hand, if they are over-identified, that is, the dimension of the moment conditions is larger than the dimension of $\bm{\alpha}$, then an $\hat{\bm{\alpha}}$ that satisfies (9) does not necessarily exist. In that case, an $\hat{\bm{\alpha}}$ satisfying
$\frac{1}{n}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}})\approx\bm{0}$
is used. Specifically, $\hat{\bm{\alpha}}$ is obtained from the generalized method of moments (Hansen 1982) or the generalized empirical likelihood method (Qin and Lawless 1994).
Here, we will use the generalized method of moments and propose $\hat{\bm{\alpha}}^{\rm CB}\equiv\mathop{\rm argmin}_{\bm{\alpha}}\{\bm{h}_{n}(\bm{\alpha})^{{\rm T}}\allowbreak\bm{W}_{n}\allowbreak\bm{h}_{n}(\bm{\alpha})\}$ as the estimator of $\bm{\alpha}$ by using the covariate balancing (CB) method, where $\bm{h}_{n}(\bm{\alpha})\equiv n^{-1}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})$ and $\bm{W}_{n}$ is an arbitrary semi-positive definite matrix. Upon defining $\bm{G}_{n}(\bm{\alpha})\equiv\partial\bm{h}_{n}(\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}$, one sees that the estimator $\hat{\bm{\alpha}}^{\rm CB}$ satisfies
$\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{h}_{n}(\hat{\bm{\alpha}}^{\rm CB})=\bm{0}.$ (10)
Its simplest candidate is the identity matrix, but the optimal matrix, $\{n^{-1}\sum_{i=1}^{n}\bm{h}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha})\allowbreak\bm{h}(\bm{d}_{i},\allowbreak\bm{x}_{i};\bm{\alpha})^{{\rm T}}\}^{-1}$, provides the most efficient estimator of $\bm{\alpha}$. However, this optimal matrix is not necessarily practical since it may be too close to being a degenerate matrix for some data, causing considerable instability in the estimation of $\bm{\alpha}$ and consequently $\bm{\theta}$.
Remark 1.
The idea of using constraints that approximately satisfy the empirical version of the moment conditions to estimate the parameters of the propensity scores has also been examined. In Ning et al. (2020b) a doubly robust estimation method is proposed for dealing with the case of high-dimensional covariates. This method combines a semiparametric estimation of the ATT with sparse estimation. It is assumed that the outcome can be expressed by a linear model of covariates, and to speed up the convergence of the Horvitz-Thompson estimator, the weak covariate balancing property, written as
$\frac{1}{n}\sum_{i=1}^{n}\bigg\{\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-1\bigg\}\bm{\alpha}^{{\rm T}}\bm{x}_{i}\approx 0,$
is included as a constraint in estimating the parameter of the propensity scores.
Here, we state the regularity conditions of the generalized method of moments (Newey and McFadden 1994) that are necessary to derive the properties of the proposed estimation method.
Assumption 5 (Consistency).
$(\mathrm{i})$ $\bm{W}_{n}$ converges in probability to $\bm{W}$, which is a semi-positive definite matrix. Also, $\bm{\alpha}=\bm{\alpha}^{\dagger}$ if and only if $\bm{W}{\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})]=\bm{0}$. $(\mathrm{ii})$ $\bm{\alpha}^{\dagger}$ is in the interior of the compact set $\bm{A}\subset\mathbb{R}^{p}$. $(\mathrm{iii})$ $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is continuous; that is, $e^{[1]}(\bm{x};\bm{\alpha})$ is continuous at each $\bm{\alpha}\in\bm{A}$ with probability one. $(\mathrm{iv})$ ${\rm E}[\sup_{\bm{\alpha}\in\bm{A}}\|\bm{h}(\bm{d},\bm{x};\bm{\alpha})\|]<\infty$.
Assumption 6 (Asymptotic normality).
$(\mathrm{i})$ $\bm{h}(\bm{d},\bm{x};\bm{\alpha})$ is continuously differentiable in a neighborhood $\Gamma$ of $\bm{\alpha}^{\dagger}$ with probability approaching one. $(\mathrm{ii})$ ${\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})]=\bm{0}$ and ${\rm E}[\|\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\|^{2}]\allowbreak<\infty$. $(\mathrm{iii})$ ${\rm E}[\sup_{\bm{\alpha}\in\Gamma}\allowbreak\|\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}\|_{\rm F}]<\infty$. $(\mathrm{iv})$ $\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\allowbreak\bm{G}(\bm{\alpha}^{\dagger})$ is non-singular for $\bm{G}(\bm{\alpha})\equiv{\rm E}[\partial\bm{h}(\bm{d},\bm{x};\bm{\alpha})/\partial\bm{\alpha}^{{\rm T}}]$.
Under these regularity conditions, the proposed method can be shown to be doubly robust, as follows. Here, we name the estimator obtained by replacing the propensity scores with $e^{[1]}(\bm{x};\hat{\bm{\alpha}}^{\rm CB})$ in (7) the covariate balancing for DID (CBD) estimator and denote it as $\hat{\bm{\theta}}^{\rm CBD}$.
Theorem 1.
If the propensity score model $e^{[1]}(\bm{x};\bm{\alpha})$ is correctly specified, or if the change in outcomes follows a linear model of covariates, that is, if
${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\bm{x}^{{\rm T}}\bm{\beta}^{[k]*}$ (11)
holds, then under Assumptions 1-6, $\hat{\bm{\theta}}^{\rm CBD}$ converges in probability to $\bm{\theta}^{*}$. That is, the proposed estimator has double robustness.
Proof.
[…]
$\sqrt{n}(\hat{\bm{\alpha}}^{\rm CB}-\bm{\alpha}^{\dagger})$
$=-\{\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\bm{G}_{n}(\hat{\bm{\alpha}}^{\rm CB})\}^{-1}\bm{G}_{n}^{{\rm T}}(\hat{\bm{\alpha}}^{\rm CB})\bm{W}_{n}\sqrt{n}\bm{h}_{n}(\bm{\alpha}^{\dagger})+{\rm o}_{\rm P}(1)$ (12)
$\stackrel{{\scriptstyle\rm d}}{{\to}}{\rm N}(\bm{0},\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\hat{\bm{\alpha}}^{\dagger})\}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{\Omega}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\hat{\bm{\alpha}}^{\dagger})\}^{-1}),$
where $\bm{\Omega}(\bm{\alpha})\equiv{\rm E}[\bm{h}(\bm{d},\bm{x};\bm{\alpha})\bm{h}(\bm{d},\bm{x};\bm{\alpha})^{{\rm T}}]$. From the consistency of $\hat{\bm{\alpha}}^{\rm CB}$, we also obtain
$\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$
$=\bigg\{\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg\}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}$ (13)
$=\bigg\{\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\bigg\}^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}+{\rm o}_{\rm P}(1).$ (14)
If the propensity score model is correctly specified and $\bm{\alpha}^{\dagger}=\bm{\alpha}^{*}$ is satisfied, that is, if $e^{[1]}(\bm{x})=e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})$ and $e^{[0]}(\bm{x})=e^{[0]}(\bm{x};\bm{\alpha}^{\dagger})$ hold, then from (5), $\hat{\bm{\theta}}^{\rm CBD}$ is a consistent estimator of $\bm{\theta}^{*}$. On the other hand, even if the propensity score model is misspecified, when the change in outcomes follows a linear model of covariates, that is, when (7) holds, we obtain
$a(\bm{x})={\rm E}[\Delta^{[1]}\mid\bm{x},d^{[1]}=1]-{\rm E}[\Delta^{[0]}\mid\bm{x},d^{[0]}=1]=\bm{x}^{{\rm T}}(\bm{\beta}^{[1]*}-\bm{\beta}^{[0]*})$
from (3). Then, it follows from (6) that
$\bm{\theta}^{*}={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[e^{[1]}(\bm{x})\bm{x}a(\bm{x})]=\bm{\beta}^{[1]*}-\bm{\beta}^{[0]*}.$
Therefore, we obtain
${\rm E}[e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\bm{x}\{\rho(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}]$
$={\rm E}\bigg[e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bigg\{\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}{\rm E}[\Delta_{i}^{[1]}\mid\bm{x}_{i},d^{[1]}_{i}=1]$
[…]
$={\rm E}\bigg[e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\bigg[\bigg\{\frac{d^{[1]}_{i}}{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\bigg\}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{1}-\bigg\{\frac{d^{[0]}_{i}}{e^{[0]}(\bm{x}_{i};\bm{\alpha}^{\dagger})}-1\bigg\}\bm{x}_{i}^{{\rm T}}\bm{\beta}^{*}_{0}\bigg]\bigg]$
$={\rm E}[\bm{H}_{1}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{1}-\bm{H}_{0}(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{\beta}^{*}_{0}]=\bm{0}.$
Here, the last equality holds from the moment conditions in (8). From the above, when the propensity score model is correctly specified, or when the change in outcomes follows a linear model of covariates, it holds that
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\bm{x}_{i}\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}={\rm O}_{\rm P}{(\sqrt{n})}$
in (14), and therefore $\hat{\bm{\theta}}^{\rm CBD}$ is consistent. ∎
Remark 2.
Let us discuss the relationship between our proposed method and the method in Sant’Anna and Zhao (2020). Unlike our setting, which estimates the conditional ATT, the latter directly estimates the ATT using a doubly robust approach. In their approach as well, the propensity scores are estimated so that the first-order moments of the distribution for the covariates are balanced. They also suppose a parametric model of the covariates for the change in the control group’s outcomes and estimate it, using these estimators to estimate the ATT. In contrast, we place importance on evaluating the heterogeneity of the ATT, including variable selection, and then set a function of the covariates for the ATT. We suppose a parametric model of the covariates for the change in outcomes as well, although not estimating it. Then, if the propensity scores are estimated so that not the first-order but the second-order moments of the distribution for the covariates are balanced, it is easily, yet somewhat unexpectedly, shown that the conditional ATT estimator is doubly robust.
3.2 Numerical experiments
Let us examine the robustness of the proposed estimation against model misspecification through simulation studies. The setting for data generation follows
$x_{1},x_{2}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+\alpha^{*}x_{2})),$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+\beta^{*}x_{1}+\epsilon^{[1]}.$
Throughout this paper, we will suppose that all random numbers are generated independently. The parameter $\beta^{*}$ represents the magnitude of the contribution from the covariate $x_{1}$ to the causal effect. The propensity score is estimated using only $x_{1}$; that is, the parameter $\alpha^{*}$ represents the magnitude of the contribution from $x_{2}$ to the true propensity scores, which is not included in the model for estimating the propensity scores. Note that when $\alpha^{*}\neq 0$, $x_{2}$ should also be included in the model, and we are doing a model misspecification because it was not included. Anyway, since (7) holds, the proposed estimation is theoretically guaranteed to be consistent. In this setting, we will compare the proposed method with a naive method estimating the propensity score by the MLE and check whether the proposed method is robust to misspecification of the propensity score model. Note that each of the numerical experiments described below was repeated 3,000 times.
The true values of the ATT, the mean estimates, and their empirical confidence intervals are listed in Table 1 for several $\alpha^{*}$, $\beta^{*}$ and sample sizes $n$. With increasing $\alpha^{*}$, which represents the magnitude of the misspecification of the propensity score model, the bias of the naive method becomes larger, and the true value is eventually no longer included in the confidence interval. Regarding the proposed method, the approximation accuracy is considerably high when the identity matrix is used as the weighting matrix in the generalized method of moments. On the other hand, when the optimal matrix is used as the weighting matrix, although the accuracy is often comparable to that of using the identity matrix, in some cases, it is similar to that of the naive method. Since there are settings in simulation studies where the optimal matrix is close to being degenerate, those cases in which the accuracy deteriorates to that of the naïve method may be attributed to the fact that the estimation of $\theta$ is not stable. Regardless of whether (7) holds, if the propensity score model is estimated correctly, both the proposed method and the naive method can estimate the ATT without bias.
Table 1: Comparison of estimates of the ATT when the propensity scores are given by the CBD method and the MLE: True is the true value of the ATT, CBD-id is the CBD method when the weighting matrix in the generalized method of moments is the identity matrix, and CBD-opt is the CBD method when it is the optimal matrix. Est indicates the average of the estimates obtained by each simulation, and Conf Int indicates its empirical 95% confidence interval.
CBD-id CBD-opt MLE
$\beta^{*}$ $\alpha^{*}$ $n$ True Est Conf Int Est Conf Int Est Conf Int
200 0.09 0.09 [-0.12, 0.30] 0.17 [-0.20, 0.51] -0.00 [-0.22, 0.21]
1.0 400 0.08 0.09 [-0.07, 0.24] 0.17 [-0.14, 0.45] -0.01 [-0.17, 0.14]
600 0.08 0.08 [-0.03, 0.21] 0.19 [-0.13, 0.12] -0.01 [-0.13, 0.12]
200 0.10 0.09 [-0.13, 0.32] 0.13 [-0.10, 0.37] -0.14 [-0.43, 0.15]
0.1 2.0 400 0.10 0.10 [-0.08, 0.25] 0.13 [-0.04, 0.29] -0.13 [-0.34, 0.06]
600 0.09 0.10 [-0.04, 0.24] 0.13 [-0.01, 0.27] -0.12 [-0.30, 0.03]
200 0.09 0.11 [-0.15, 0.39] 0.12 [-0.14, 0.37] -0.31 [-0.80, 0.10]
3.0 400 0.09 0.11 [-0.07, 0.29] 0.13 [-0.06, 0.31] -0.30 [-0.60, -0.03]
600 0.10 0.11 [-0.03, 0.26] 0.12 [-0.02, 0.27] -0.30 [-0.55, -0.08]
1.0 200 0.42 0.43 [0.21, 0.65] 0.52 [0.09, 0.93] 0.31 [0.06, 0.57]
600 0.43 0.43 [0.31, 0.56] 0.55 [0.20, 0.80] 0.31 [0.18, 0.45]
0.5 2.0 200 0.47 0.46 [0.23, 0.70] 0.50 [0.27, 0.75] 0.21 [-0.10, 0.51]
600 0.46 0.47 [0.33, 0.61] 0.51 [0.37, 0.64] 0.22 [0.04, 0.39]
3.0 200 0.47 0.50 [0.24, 0.78] 0.51 [0.25, 0.77] 0.05 [-0.45, 0.47]
600 0.48 0.50 [0.35, 0.66] 0.51 [0.37, 0.66] 0.05 [-0.20, 0.28]
1.0 200 0.85 0.85 [0.62, 1.10] 0.97 [0.47, 1.44] 0.71 [0.44, 1.01]
600 0.85 0.85 [0.72, 0.99] 1.00 [0.57, 1.31] 0.71 [0.55, 0.87]
1.0 2.0 200 0.92 0.93 [0.68, 1.19] 0.97 [0.73, 1.24] 0.63 [0.30, 0.96]
600 0.92 0.93 [0.78, 1.08] 0.98 [0.83, 1.12] 0.65 [0.46, 0.83]
3.0 200 0.95 0.99 [0.71, 1.28] 1.00 [0.73, 1.27] 0.50 [-0.02, 0.93]
600 0.95 0.99 [0.83, 1.15] 1.00 [0.85, 1.16] 0.50 [0.24, 0.74]
200 2.56 2.55 [2.13, 2.96] 2.77 [1.84, 3.49] 2.30 [1.82, 2.82]
1.0 400 2.56 2.55 [2.27, 2.86] 2.78 [1.96, 2.67] 2.30 [1.95, 2.67]
600 2.56 2.55 [2.33, 2.76] 2.82 [2.03, 3.32] 2.30 [2.02, 2.58]
200 2.75 2.78 [2.42, 3.18] 2.86 [2.51, 3.24] 2.35 [1.91, 2.86]
3.0 2.0 400 2.75 2.78 [2.51, 3.04] 2.85 [2.60, 3.11] 2.35 [2.01, 2.68]
600 2.75 2.78 [2.57, 3.00] 2.85 [2.66, 3.06] 2.35 [2.07, 2.63]
200 2.85 2.93 [2.55, 3.33] 2.95 [2.59, 3.32] 2.28 [1.70, 2.85]
3.0 400 2.85 2.93 [2.67, 3.21] 2.95 [2.70, 3.21] 2.28 [1.89, 2.67]
600 2.85 2.93 [2.71, 3.15] 2.95 [2.75, 3.16] 2.28 [1.97, 2.59]
4 Model selection criterion for semiparametric difference-in-differences when propensity scores are known
4.1 Risk function and its asymptotic evaluation in semiparametric difference-in-differences
As mentioned in Section 1, there are no reasonable model selection criteria even for the basic setting treated in Abadie (2005). To remedy this situation, in this section, we will first derive a model selection criterion for when the propensity scores are known and then develop a model selection criterion for when the proposed CBD method is used. Specifically, by using the ideas of Baba et al. (2017) and Baba and Ninomiya (2025), we will derive the model selection criteria as asymptotically unbiased estimators of the natural risk corresponding to the loss function used to obtain SDID estimators. Using these criteria, the dimension of $\bm{\theta}$ is determined and the component of $\bm{x}$ is selected. A natural choice is to define a weighted risk based on (4) as
$\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{{\rm E}[y^{[1]}_{i}(1)-y^{[0]}_{i}(1)\mid\bm{x}_{i},d^{[1]}_{i}=1]-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}]$
$=\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}]$
$\ \phantom{=}-\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-{\rm E}[y^{[1]}_{i}(1)-y^{[0]}_{i}(1)\mid\bm{x}_{i},d^{[1]}_{i}=1]\}^{2}]$
$\ \phantom{=}+2\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-{\rm E}[y^{[1]}_{i}(1)-y^{[0]}_{i}(1)\mid\bm{x}_{i},d^{[1]}_{i}=1]\}$
[…]
Following the derivation of the conventional $C_{p}$ criterion, we will remove the expectation in the first term, ignore the second term which is independent of the models, and asymptotically evaluate the third term. For simplicity, we will refer to the third term as a bias when the risk is simply estimated. If we define $\varepsilon_{i}\equiv\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-a(\bm{x}_{i})$ using $a(\bm{x}_{i})$ in (1), the third term can be written as
$2\sum_{i=1}^{n}{\rm E}[e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-a(\bm{x}_{i})\}\{\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}-a(\bm{x}_{i})\}]$
$=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}-a(\bm{x}_{i})\}]=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\bm{x}_{i}^{{\rm T}}(\hat{\bm{\theta}}-\bm{\theta}^{*})].$ (16)
The second equality is satisfied by ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})a(\bm{x}_{i})]=0$ and ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\allowbreak\bm{x}^{{\rm T}}_{i}\bm{\theta}^{*}]=0$ since it follows from (3) that ${\rm E}[\varepsilon_{i}\mid\bm{x}_{i}]=0$. Here, we have used the same $e^{[1]}(\bm{x}_{i};\bm{\alpha})$ as that was used to construct $\hat{\bm{\theta}}$. Then, letting $c^{\rm limit}$ be the weak limit of the third term with the expectation removed, we consider ${\rm E}[c^{\rm limit}]$ as the asymptotic evaluation of the third term of (15). Note that while we consider a risk based on the loss function in deriving the SDID estimator, it is also possible to consider a simple squared error.
4.2 Derivation of model selection criterion
In the case of known propensity scores, we will denote $e^{[1]}(\bm{x}_{i};\bm{\alpha}^{*})$ as $e^{[1]}(\bm{x}_{i})$ and $\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha}^{*})$ as $\rho(\bm{t}_{i},\bm{x}_{i})$ for simplicity. The main term of the error of the SDID estimator can be expressed as
$\hat{\bm{\theta}}-\bm{\theta}^{*}={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\{\rho(\bm{t}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\{1+{\rm o}_{\rm P}(1)\}$
$={\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\{\varepsilon_{i}+a(\bm{x}_{i})-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\{1+{\rm o}_{\rm P}(1)\}.$
Accordingly, (16) is evaluated as
${\rm E}[c^{\rm limit}]=2\sum_{i=1}^{n}{\rm E}\bigg[\varepsilon_{i}e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}\frac{1}{n}\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j})\bm{x}_{j}\{\varepsilon_{j}+a(\bm{x}_{j})-\bm{x}_{j}^{{\rm T}}\bm{\theta}^{*}\}\bigg].$
This expectation for the case of $i\neq j$ is the product of the expectations for $i$ and $j$, from the independence among the samples. The expectation for $i$ can be calculated as
${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}]={\rm E}_{\bm{x}_{i}}[{\rm E}[\varepsilon_{i}\mid\bm{x}_{i}]e^{[1]}(\bm{x}_{i})\bm{x}_{i}^{{\rm T}}]=\bm{0}.$ (17)
Therefore, we only have to consider the case of $i=j$. Since $a(\bm{x}_{j})-\bm{x}_{j}^{{\rm T}}\bm{\theta}^{*}$ contains only $\bm{x}_{j}$ as a random variable, this term disappears from (17), and the following theorem is obtained.
Theorem 2.
When the propensity scores are known, it holds that
${\rm E}[c^{\rm limit}]=2{\rm E}[\varepsilon e^{[1]}(\bm{x})\bm{x}^{{\rm T}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}e^{[1]}(\bm{x})\bm{x}\varepsilon]$
$=2{\rm tr}\{{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[\sigma^{2}(\bm{x})e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]\},$
where
$\sigma^{2}(\bm{x})\equiv{\rm E}[\varepsilon^{2}\mid\bm{x}]$ $={\rm E}[\{\rho(\bm{t},\bm{x})\Delta\}^{2}\mid\bm{x}]-{\rm E}[\rho(\bm{t},\bm{x})\Delta\mid\bm{x}]^{2}.$ (18)
Although one may think that an empirical evaluation of ${\rm E}[c^{\rm limit}]$ is easy, we must consider how to deal with the term ${\rm E}[{\rm E}[\rho(\bm{t},\bm{x})\Delta\mid\bm{x}]^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]$. This cannot be empirically evaluated as $n^{-1}\sum_{i=1}^{n}\rho(\bm{t},\bm{x})^{2}\allowbreak\Delta^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}$. Here, we will substitute $\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}$ into ${\rm E}[\rho(\bm{t}_{i},\bm{x}_{i})\Delta\mid\bm{x}_{i}]$. As a result, when the propensity scores are known, we propose
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\{\rho(\bm{t}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}$
$+2{\rm tr}\biggl(\biggl\{\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr\}^{-1}\biggl[\sum_{i=1}^{n}\{\rho(\bm{t}_{i},\bm{x}_{i})^{2}\Delta^{2}-(\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}})^{2}\}e^{[1]}(\bm{x}_{i})^{2}\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr]\biggr)$ (19)
as the model selection criterion for the basic SDID approach.
4.3 Extension of QICw
As a comparator for the proposed model selection criterion, we will extend QICw (Platt et al. 2013), which is an intuitive model selection criterion for weighted estimation using propensity scores, to the setting of this paper. First, we construct the goodness-of-fit term using the same idea as for QICw. To do so, the log-likelihood function for full data is differentiated, a weighted potential outcome variable is substituted into the one whose expectation is the same as the outcome variable, the differentiation is reverted, and the estimator is substituted into the parameter. The number of parameters is used as the penalty term. Within this setting, the log-likelihood function for full data is
$-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}[\{\mbox{$y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$}\}-\bm{x}_{i}^{{\rm T}}\bm{\theta}]^{2}.$
If we take $\rho_{i}\Delta_{i}$ as the weighted potential outcome variable, its conditional expectation can be shown to be the same as the conditional expectation of the outcome variable from (3). We will call the sum of the goodness-of-fit term and the penalty term multiplied by a constant $2\sigma^{2}$, that is,
$\sum_{i=1}^{n}\{\rho(\bm{d}_{i},\bm{x}_{i})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}\}^{2}+2\sigma^{2}p,$
the QICw for the SDID approach. In fact, if $\sigma^{2}(\bm{x})$ is a constant, then our proposed criterion is equivalent to the QICw. The problem is to estimate the variance $\sigma^{2}$ of the difference, $y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$. This estimate cannot be obtained from the data using only the present assumptions. Since it follows from $y^{[0]}(0)=y^{[1]}(0)$ that $y^{[1]}_{i}(1)-y^{[0]}_{i}(1)=\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$, and from Assumption 1, the expectation of $\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$ is the same as the expectation of $\Delta_{i}^{[1]}$ in $d^{[0]}_{i}=0$ minus $\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$. However, in order to evaluate the variance, we will assume that $\Delta_{i}^{[1]}$ and $\Delta_{i}^{[0]}$ are independent conditioned on $d^{[1]}=1$ and $d^{[0]}=1$, from which we have
$\sigma^{2}={\rm Var}[\Delta^{[1]}\mid d^{[1]}=1]+{\rm Var}[\Delta^{[0]}\mid d^{[0]}=1]$
$={\rm E}[\Delta^{2}\mid d^{[1]}=1]-{\rm E}[\Delta\mid d^{[1]}=1]^{2}+{\rm E}[\Delta^{2}\mid d^{[0]}=1]-{\rm E}[\Delta\mid d^{[0]}=1]^{2}.$
Moreover, defining $n_{1}\equiv\sum_{i=1}^{n}d^{[1]}_{i}$ and $n_{0}\equiv\sum_{i=1}^{n}d^{[0]}_{i}$, we will use
$\hat{\sigma}^{2}=\frac{1}{n_{1}}\sum_{i=1}^{n}d^{[1]}_{i}\Delta_{i}^{2}-\bigg(\frac{1}{n_{1}}\sum_{i=1}^{n}d^{[1]}_{i}\Delta_{i}\bigg)^{2}+\frac{1}{n_{0}}\sum_{i=1}^{n}d^{[0]}_{i}\Delta_{i}^{2}-\bigg(\frac{1}{n_{0}}\sum_{i=1}^{n}d^{[0]}_{i}\Delta_{i}\bigg)^{2}$
in the present study.
Remark 3.
Let us compare $\sigma^{2}(\bm{x})$ in the proposed criterion with $\sigma^{2}$ in the QICw. In (18), if all expectations do not actually depend on $\bm{x}$, then
$\sigma^{2}(\bm{x})=\$${\rm E}\bigg[\frac{\Delta^{[1]2}}{e^{[1]}}\ \bigg|\ d^{[1]}=1\bigg]-{\rm E}[\Delta^{[1]}\mid d^{[1]}=1]^{2}+{\rm E}\bigg[\frac{\Delta^{[0]2}}{e^{[0]}}\ \bigg|\ d^{[0]}=1\bigg]-{\rm E}[\Delta^{[0]}\mid d^{[0]}=1]^{2}$
$+2{\rm E}[\Delta^{[1]}\mid d^{[1]}=1]{\rm E}[\Delta^{[0]}\mid d^{[0]}=1].$
Since $\Delta^{[1]}$ and $\Delta^{[0]}$ include fluctuations over time, it is natural that $\Delta^{[1]}\Delta^{[0]}$ should be non-negative. In other words, basically $\sigma^{2}(\bm{x})>\sigma^{2}$, and the penalty term of the proposed criterion is larger than that of the QICw.
4.4 Numerical experiments
Here, the performance of the proposed criterion (19) when the propensity scores are known is evaluated in simulation studies. The data generation settings include Case 1-1,
$x_{1}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1})),$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+1+\beta^{*}x_{1}+\epsilon^{[1]},$
and Case 1-2,
$x_{1},x_{2}\sim{\rm Uniform}(0,2),\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})),$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+1+\beta^{*}(x_{1}+x_{2})+\epsilon^{[1]}.$
In both cases, $\beta^{*}$ represents the magnitude of contribution of the covariate $x_{1}$ to the causal effect.
Table 2 shows whether the second term in (19), the penalty term of the proposed criterion, accurately approximates the bias in (16) that we originally wanted to evaluate. For all values of $\beta^{*}$ and sample size $n$, our penalty terms are close to the true values of the bias, while the evaluated penalty terms in the QICw significantly underestimate the bias.
Table 2: Bias evaluation for the case where the propensity scores are known: True denotes the true value of the bias evaluated by the Monte Carlo method, and Proposal the evaluated penalty term of the proposed criterion. The evaluated penalty term of the QICw is listed as a reference.
Case 1-1 Case 1-2
$\beta^{*}$ $n$ True Proposal QICw True Proposal QICw
200  37.54  37.29 2.23  39.20  39.43  5.96
0.1 400  38.56  37.71 2.24  39.75  39.74  5.98
600  37.89  37.89 2.27  39.79  39.68  6.00
200  57.34  56.28 2.31  59.77  58.69  6.39
0.5 400  59.36  56.28 2.33  59.36  59.52  6.48
600  58.56  57.35 2.34  60.20  59.37  6.48
200  92.48  91.33 2.56 103.24  98.85  7.91
1.0 400  98.17  91.43 2.58 108.09  99.46  7.99
600  87.03  92.62 2.59  98.33  99.57  7.99
200 359.64 358.56 5.17 439.35 429.91 23.65
3.0 400 345.73 348.96 5.20 441.47 430.88 24.05
600 367.73 365.75 5.22 435.32 430.09 23.87
Next, let us examine the performance of the model selection using the proposed criterion, with data generated from
$x_{1},x_{2},\ldots,x_{l}\sim{\rm Uniform}(0,2),\quad\bm{x}=(1,x_{1},x_{2},\ldots,x_{l})^{{\rm T}},$
$y^{[0]}(0)=y^{[1]}(0)\sim{\rm N}(0,1),\quad\epsilon^{[0]},\epsilon^{[1]}\sim{\rm N}(0,1),$
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+\bm{x}^{{\rm T}}\bm{\beta}^{[1]*}+\epsilon^{[1]},$
in Case 2-1,
$l=4,\quad\bm{\beta}^{[1]*}=(1,\beta^{*},\bm{0}_{3}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1})),$
Case 2-2,
$l=4,\quad\bm{\beta}^{[1]*}=(1,\bm{1}_{2}^{{\rm T}}\beta^{*},\bm{0}_{2}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})),$
and Case 2-3,
$l=6,\quad\bm{\beta}^{[1]*}=(1,\bm{1}_{2}^{{\rm T}}\beta^{*},\bm{0}_{4}^{{\rm T}})^{{\rm T}},\quad d^{[1]}\sim{\rm Bernoulli}({\rm logit}(-x_{1}+x_{2})).$
Table 3 compares the proposed criterion and the QICw when using the forward selection method. As the main index to measure the goodness of the criteria, we use an empirical estimate of the risk, specifically, the average of 3,000 values of
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\bm{\alpha})(\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}})^{2}.$
In Case 2-2, where there are relatively few unnecessary covariates among the candidates, that is, where the negative effect of false positives is less likely to appear, the QICw slightly outperforms the proposed criterion for $\beta^{*}=0.5$ and $\beta^{*}=1.0$. However, the proposed criterion outperforms the QICw for the other values of beta in Case 2-2, and it outperforms the QICw for all values in Case 2-1 and Case 2-3, i.e., settings with many covariates are unrelated to the true structure. Here, since the QICw underestimates the penalty term, the false positive must be large and too many covariates are selected in all settings.
Table 3: Comparison of the proposed criterion and the QICw through model selection for case in which the propensity scores are known: Risk denotes the empirical estimate of the risk of the selected model, TP the average number of covariates that are true positive, and FP the average number of covariates that are false positive.
Case 2-1 Case 2-2 Case 2-3
$\beta^{*}$ $n$ Risk TP FP Risk TP FP Risk TP FP
200 Proposal 11.05 0.50 1.29  14.07 0.61 0.62  19.01 0.63 1.15
QICw 11.29 0.53 1.47  15.94 0.90 0.95  21.74 0.90 1.73
0.1 400 Proposal 11.04 0.47 1.30  13.99 0.68 0.52  19.72 0.69 1.08
QICw 11.27 0.51 1.49  15.23 0.98 0.86  22.45 0.98 1.77
600 Proposal 10.86 0.48 1.25  14.40 0.75 0.51  19.51 0.72 1.14
QICw 11.00 0.51 1.44  15.88 1.07 0.83  22.62 1.02 1.81
200 Proposal 14.68 0.70 1.35  23.20 1.23 0.61  30.01 1.20 1.21
QICw 14.94 0.74 1.61  23.33 1.50 1.00  31.61 1.48 1.96
0.5 400 Proposal 14.42 0.83 1.29  22.95 1.60 0.60  31.82 1.55 1.14
QICw 14.71 0.86 1.58  22.48 1.78 1.01  32.67 1.75 1.93
600 Proposal 14.67 0.89 1.31  23.24 1.74 0.57  30.25 1.75 1.20
QICw 14.93 0.91 1.61  23.00 1.88 0.98  31.96 1.88 1.98
200 Proposal 21.69 0.89 1.39  37.19 1.69 0.61  48.24 1.69 1.26
QICw 22.00 0.91 1.67  36.59 1.85 1.04  49.53 1.84 2.18
1.0 400 Proposal 21.09 0.97 1.40  34.65 1.95 0.63  46.07 1.94 1.20
QICw 21.61 0.98 1.76  35.71 1.99 1.08  49.45 1.98 2.12
600 Proposal 20.74 1.00 1.35  32.16 1.99 0.60  43.66 2.00 1.24
QICw 21.30 1.00 1.72  34.38 1.99 1.07  48.02 2.00 2.19
200 Proposal 67.79 0.99 1.54 136.61 1.96 0.67 180.41 1.96 1.33
QICw 69.22 1.00 1.93 141.09 1.99 1.19 194.02 1.98 2.34
3.0 400 Proposal 66.20 1.00 1.57 133.32 2.00 1.20 175.10 2.00 1.25
QICw 67.83 1.00 1.97 141.74 2.00 1.20 191.79 2.00 2.38
600 Proposal 65.76 1.00 1.48 126.36 2.00 0.62 173.78 2.00 1.29
QICw 67.51 1.00 1.91 135.24 2.00 1.17 191.42 2.00 2.36
5 Model selection criterion for covariate balancing difference-in-differences
5.1 Derivation of model selection criterion
As we did for the case of known propensity scores in Section 4.2, here, we will derive a validated model selection criterion for when the CBD method is used. Let us evaluate
$\sum_{i=1}^{n}2{\rm E}[e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{t}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-{\rm E}[y^{[1]}_{i}-y^{[0]}_{i}\mid\bm{x}_{i},d^{[1]}_{i}=1]\}\bm{x}_{i}^{{\rm T}}(\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*})],$
which is the third term of the risk in (16). Substituting $\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}$ into ${\rm E}[y^{[1]}_{i}-y^{[0]}_{i}\mid\bm{x}_{i},d^{[1]}_{i}=1]$ first and then substituting the right-hand side of (13) into $\hat{\bm{\theta}}^{\rm CBD}-\bm{\theta}^{*}$ yields
$2{\rm E}\bigg[\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{t}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}^{{\rm T}}\bigg\{\frac{1}{n}\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{j}\bm{x}_{j}^{{\rm T}}\bigg\}^{-1}$
[…]
Since $n^{-1}\sum_{j=1}^{n}e^{[1]}(\bm{x}_{j};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{j}\bm{x}_{j}^{{\rm T}}$ converges in probability to
$\bm{L}(\bm{\alpha}^{\dagger})\equiv{\rm E}[e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\bm{x}\bm{x}^{{\rm T}}],$ (20)
by letting $\bm{u}$ be the Gaussian random vector that is the weak limit of
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i},$ (21)
we can express the weak limit of (16) with expectation removed as
$c^{\rm limit}=2\bm{u}^{{\rm T}}\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{u}=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{u}\bm{u}^{{\rm T}}\}.$
When $\hat{\bm{\theta}}^{\rm CBD}$ is a consistent estimator of $\bm{\theta}^{*}$, it holds that ${\rm E}[\bm{u}\bm{u}^{{\rm T}}]={\rm Var}[\bm{u}]$ since the expectation of $\bm{u}$ becomes $\bm{0}$. Accordingly, the asymptotic evaluation of (16), which is to be used in the information criterion, can be written as
${\rm E}[c^{\rm limit}]=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}{\rm Var}[\bm{u}]\}.$
Now let us derive a specific expression for ${\rm Var}[\bm{u}]$. Since the error of the estimator $\hat{\bm{\alpha}}^{\rm CB}$ in the propensity scores is written as (12), we can expand (21) as
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}$
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\bm{x}_{i}\bigg\{e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}$
[…]
$=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}[e^{[1]}(\bm{x}_{i};\bm{\alpha}^{\dagger})\{\rho(\bm{d}_{i},\bm{x}_{i};\bm{\alpha}^{\dagger})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}_{i}$
[…]
where
$\bm{M}(\bm{\alpha},\bm{\theta})={\rm E}\biggl[\dfrac{\partial}{\partial\bm{\alpha}}e^{[1]}(\bm{x};\bm{\alpha})\biggl[\frac{d^{[1]}-1}{\{1-e^{[1]}(\bm{x};\bm{\alpha})\}^{2}}\Delta-\bm{x}^{{\rm T}}\bm{\theta}\biggr]\bm{x}\biggr].$
Then, by defining
$\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\equiv{\rm Var}$ $[e^{[1]}(\bm{x};\bm{\alpha}^{\dagger})\{\rho(\bm{d},\bm{x};\bm{\alpha}^{\dagger})\Delta-\bm{x}^{{\rm T}}\bm{\theta}^{*}\}\bm{x}$
$-\bm{M}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\{\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{G}(\bm{\alpha}^{\dagger})\}^{-1}\bm{G}^{{\rm T}}(\bm{\alpha}^{\dagger})\bm{W}\bm{h}(\bm{d},\bm{x};\bm{\alpha}^{\dagger})],$ (22)
it can be shown that ${\rm Var}[\bm{u}]=\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$. Thus, the following theorem is obtained.
Theorem 3.
When the propensity scores are estimated using the CBD method, the third term in (15) is evaluated as
${\rm E}[c^{\rm limit}]=2{\rm tr}\{\bm{L}(\bm{\alpha}^{\dagger})^{-1}\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})\},$
where $\bm{L}(\cdot)$ and $\bm{V}(\cdot,\cdot)$ are defined in (20) and (22), respectively.
Since the penalty term in this theorem depends on the true values of the parameters, we can use empirical estimates in practice. Specifically, we will replace $\bm{L}(\bm{\alpha}^{\dagger})$ by $\bm{L}_{n}(\hat{\bm{\alpha}}^{\rm CB})\equiv n^{-1}\sum_{i=1}^{n}\allowbreak e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}$ and $\bm{V}(\bm{\alpha}^{\dagger},\bm{\theta}^{*})$ by
$\bm{V}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})$
$\equiv\frac{1}{n}\sum_{i=1}^{n}[e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD}\}\bm{x}_{i}$
[…]
where
$\bm{M}_{n}(\bm{\alpha},\bm{\theta})\equiv\frac{1}{n}\sum_{i=1}^{n}\dfrac{\partial}{\partial\bm{\alpha}}e^{[1]}(\bm{x}_{i};\bm{\alpha})\bigg[\frac{d^{[1]}_{i}-1}{\{1-e^{[1]}(\bm{x}_{i};\bm{\alpha})\}^{2}}\Delta_{i}-\bm{x}_{i}^{{\rm T}}\bm{\theta}\bigg]\bm{x}_{i}.$ (23)
As a result, for when the propensity scores are estimated with the CBD method, we propose
$\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\{\rho(\bm{d}_{i},\bm{x}_{i};\hat{\bm{\alpha}}^{\rm CB})\Delta_{i}-\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}^{\rm CBD}\}^{2}+2{\rm tr}\{\bm{L}_{n}(\hat{\bm{\alpha}}^{\rm CB})^{-1}\bm{V}_{n}(\hat{\bm{\alpha}}^{\rm CB},\hat{\bm{\theta}}^{\rm CBD})\}$ (24)
as the model selection criterion, which is an asymptotically unbiased estimator of the risk in (15). Here, the propensity scores are estimated using covariate balancing, but the information criterion can be derived in the same way when the MLE is used (see Supplementary Material).
5.2 Numerical experiments
The simulation studies presented in this section examine the accuracy of the approximation for the penalty term and the performance of the model selection criterion in (24) for the proposed CBD method. On the basis of the results in Section 4.4, we will use the identity matrix as the weighting matrix in the generalized method of moments and note that, even when the optimal matrix is used as the weighting matrix, the accuracy of the approximation for the penalty term and the performance of the model selection with comparable to those obtained with the identity matrix (see the Supplementary Material).
Table 4 lists results for data generated as in Cases 1-1 and Cases 1-2 in Section 4.4 to check whether the second term in (24), the penalty term of the proposed criterion, accurately approximates the bias in (16), which we originally wanted to evaluate, for various values of the parameter $\beta^{*}$ and sample size $n$. In all cases, the evaluated penalty terms of the proposed criterion are much closer to the bias than the evaluated penalty terms of the QICw, and its approximation accuracy is high. In contrast, the penalty term of the QICw significantly underestimates the bias in all cases.
Table 4: Bias evaluation for case in which the propensity scores are estimated using the CBD method with the identity matrix as the weighting matrix: True denotes the true value of the bias evaluated by the Monte Carlo method, and Proposal the penalty term of the proposed criterion. The penalty term of the QICw is also listed for reference.
Case 1-1 Case 1-2
$\beta^{*}$ $n$ True Proposal QICw True Proposal QICw
200  7.53  7.35 2.23 17.61 16.98  5.94
0.1 400  7.19  7.31 2.24 17.88 17.55  5.99
600  7.09  7.34 2.27 17.10 17.35  5.98
200  8.21  8.04 2.31 19.42 18.76  6.38
0.5 400  7.79  7.89 2.33 19.51 19.62  6.48
600  7.75  7.88 2.34 18.94 19.25  6.47
200  9.47  9.56 2.58 23.20 22.67  7.92
1.0 400  9.04  9.25 2.58 23.07 23.18  8.00
600  8.96  9.14 2.59 22.62 22.98  8.00
200 19.10 19.22 5.17 54.55 54.82 23.62
3.0 400 18.36 18.58 5.20 53.71 54.81 24.08
600 17.88 18.49 5.22 53.55 54.64 23.90
Table 5 lists results for data generated as in Case 2-1, Case 2-2, and Case 2-3 in Section 4.4 and comparing the proposed criterion and the QICw through the forward selection method. The main index of goodness for the model selection criteria is again the empirical estimate of the risk. Unlike Section 4.4, where the propensity scores are known, the proposed criterion always outperforms the QICw in all cases, even in Case 2-2, where relatively few covariates are unnecessary among the candidates. The proposed criterion is especially superior when the candidates include many covariates that do not contribute to the true structure, as in Case 2-1 and Case 2-3 of Section 4.4. The QICw underestimates the penalty term; thus, too many covariates are selected.
Table 5: Comparison of the proposed criterion and the QICw through model selection for case in which the propensity scores are estimated using the CBD method with the identity matrix as the weighting matrix: Risk denotes the empirical estimate of the risk of the selected model, TP the average number of covariates that are true positive, and FP the average number of covariates that are false positive.
Case 2-1 Case 2-2 Case 2-3
$\beta^{*}$ $n$ Risk TP FP Risk TP FP Risk TP FP
200 Proposal 7.33 0.21 0.53 12.38 0.41 0.33 16.49 0.38 0.74
QICw 10.32 0.48 1.55 16.96 0.91 0.92 22.57 0.84 1.79
0.1 400 Proposal 7.30 0.21 0.49 12.06 0.43 0.33 17.02 0.46 0.69
QICw 10.13 0.47 1.46 15.75 0.96 0.91 22.47 0.96 1.74
600 Proposal 6.88 0.20 0.46 12.01 0.46 0.34 16.62 0.50 0.66
QICw 9.57 0.50 1.43 15.11 0.96 0.91 21.91 0.98 1.79
200 Proposal 10.73 0.51 0.55 20.12 1.12 0.34 26.11 1.11 0.72
QICw 13.12 0.77 1.65 21.56 1.55 1.01 30.24 1.51 2.01
0.5 400 Proposal 10.64 0.72 0.50 20.34 1.52 0.33 26.36 1.54 0.69
QICw 12.87 0.88 1.61 20.47 1.80 1.02 29.42 1.80 1.97
600 Proposal 10.02 0.84 0.48 17.98 1.79 0.33 24.29 1.77 0.66.
QICw 12.05 0.97 1.60 19.15 1.92 1.03 28.46 1.92 2.01
200 Proposal 13.83 0.86 0.54 25.04 1.78 0.33 34.29 1.80 0.72
QICw 17.55 0.95 1.78 28.00 1.92 1.10 41.46 1.93 2.15
1.0 400 Proposal 12.44 0.98 0.50 22.03 1.97 0.33 31.09 1.97 0.67
QICw 16.82 1.00 1.76 26.42 2.00 1.10 40.19 2.00 2.16
600 Proposal 11.33 1.00 0.48 20.06 2.00 0.31 28.79 2.00 0.63
QICw 15.91 1.00 1.76 25.28 2.00 1.11 39.30 2.00 2.17
200 Proposal 34.16 1.00 0.52 60.56 2.00 0.31 95.78 2.00 0.68
QICw 48.77 1.00 1.94 81.22 2.00 1.17 136.38 2.00 2.37
3.0 400 Proposal 31.96 1.00 0.51 59.75 2.00 0.31 91.38 2.00 0.65
QICw 46.25 1.00 1.95 80.71 2.00 1.18 132.43 2.00 2.36
600 Proposal 29.56 1.00 0.49 57.25 2.00 0.31 87.98 2.00 0.61
QICw 44.02 1.00 1.95 78.45 2.00 1.18 130.44 2.00 2.34
6 Real data analysis
The LaLonde dataset, originally presented in LaLonde (1986), is included in the R package Matching version 4.10.14. In the dataset, the group that took the U.S. job training program in 1976 is denoted as $t=1$ and the group that did not take the program is denoted as $t=0$. The difference in annual income in 1978 after the training for each group is estimated via the ATT. Age (age, years of education (educ), income in 1974 (re74), black (black), hispanic (hisp), married (married), and high school graduate or higher (nodegr) are covariates, i.e. $p=7$, and the outcome variable is the difference between income in 1978 (re78) and income in 1974 (re78). The sample size is $n=445$. The objective of our experiment here is not to extract information from these data, but to determine whether the proposed method makes a significant difference from an existing method. Therefore, to reduce the sample size of the dataset, we divided the data into three parts and analyzed each of them. Specifically, for every integer $k$, the 3k23k-2-th row of data are included in Block 1, the 3k13k-1-th row in Block 2, and the $3k$-th row in Block 3, respectively.
First, we estimate the propensity scores by using the covariate balancing with the identity matrix as the weighting matrix. Then, supposing that the ATT can be written as a linear sum of covariates, we estimate their coefficients by using the SDID approach. In Table 5, the covariates are selected using the proposed criterion or the QICw with the forward selection method and the coefficients of the selected covariates are estimated. The QICw selects all covariates in any block, while the proposed criterion does not select some variables. In particular, the difference between the models selected for Block 1 and Block 3 is rather large. Because we do not know the true structure of the real data, it is impossible to judge which of the two criteria is best; however, we can see that the difference between them is quite large. These results indicates that, in a setting such as here, it is essential to use a model selection criterion that is theoretically valid rather than an intuitive one.
Table 6: Comparison of the proposed criterion and the QICw through model selection for the LaLonde dataset. The numbers with exponential notation are the estimated coefficients for each covariate, and “0” means that the covariate is not selected in the model selection.
Block intercept age educ re74
1 Proposal -2.14E+3 0E+0 0E+0 0E+0
QICw -1.37E+4 5.29E+2 4.43E+1 -2.01E-1
2 Proposal 1.59E+4 0E+0 1.08E+3 2.54E-1
QICw 2.38E+4 -1.91E+2 8.22E+2 2.73E-1
3 Proposal -8.50E+2 0E+0 0E+0 0E+0
QICw -8.51E+3 6.13E+0 3.35E+2 2.56E-1
Block black hisp married nodegr
1 Proposal 0E+0 0E+0 0E+0 0E+0
QICw 3.55E+2 -8.19E+3 -9.41E+2 -1.30E+3
2 Proposal -1.86E+4 -2.34E+4 4.02E+3 -1.26E+4
QICw -1.85E+4 -2.42E+4 4.88E+3 -1.34E+4
3 Proposal 0E+0 -1.03E+4 0E+0 0E+0
QICw 3.84E+3 -6.71E+3 2.92E+3 -5.90E+2
7 Discussion
Our proposal, the CBD method, is an SDID approach incorporating covariate balancing. In order to ensure double robustness, i.e. to ensure that the ATT is consistently estimated when the model for the propensity score is correctly specified or the model for the changes in the outcomes over time is correctly supposed, we have shown that the second-order moments of the distribution for the covariates should be balanced between the treatment and control groups with the former as the target population. Since covariate balancing usually balances the first-order moments, the value for the fact that the moment is second-order rather than first-order lies more in having been discovered due to its unexpectedness than in how to demonstrate its validity. Moreover, our model selection criterion for the CBD method is an asymptotically unbiased estimator of risk based on the loss function used for the estimation. Specifically, we derived a model selection criterion using just a DID-specific assumption called the conditional parallel trend assumption, i.e., without relying on the ignorable treatment assignment condition, which is often assumed in causal inference. The penalty term has been shown to be considerably different from twice the number of parameters that often appears in the AIC-type model selection criteria. Since no model selection criteria existed for even the basic SDID approach, we developed a model selection criterion not only for the CBD method, but also for when the propensity scores are known or estimated by MLE. Numerical experiments show that the proposed method estimates the ATT more robustly than the method that estimates propensity scores by using MLE and that the proposed criterion clearly reduces the risk targeted by the SDID approach compared with the intuitive generalization of the existing information criterion to the SDID approach. In addition, analyses of real data confirm that there is a large difference between the results of the proposed method and those of the existing method.
Although this paper deals with a simple linear model of covariates as a working model for the outcome, its results can be easily extended to situations where a nonlinear model of covariates should be used, for example, when using the kernel method. In fact, Hazlett (2020) has treated the function $\bm{\phi}(\bm{x}):\mathbb{R}^{p}\rightarrow\mathbb{R}^{q}$ in a reproducing kernel Hilbert space where $q$ may be larger than $p$ or $n$. They assumed that the outcome could be represented by a linear model of $\bm{\phi}(\bm{x})$, which in our setting can be written as
${\rm E}[y^{[1]}(1)-y^{[0]}(1)\mid\bm{x},d^{[1]}=1]=\bm{\phi}(\bm{x})^{{\rm T}}\bm{\theta}.$ (25)
Then, by using
$\sum_{i=1}^{n}\bigg\{\frac{d^{[1]}_{i}\bm{\phi}(\bm{x}_{i})}{e^{[1]}(\bm{x}_{i};\bm{\alpha})}-\frac{d^{[0]}_{i}\bm{\phi}(\bm{x}_{i})}{e^{[0]}(\bm{x}_{i};\bm{\alpha})}\bigg\}=\bm{0}$
as the moment conditions for estimating the propensity scores, they propose the kernel balancing method, which can flexibly estimate causal effects even if the relationship between the covariate and the outcome is nonlinear. For making a doubly robust estimation using the kernel balancing in the setting of this paper, together with the assumption in (25), let us additionally suppose that the change in the outcomes follows
${\rm E}[\Delta^{[k]}\mid\bm{x},d^{[k]}=1]=\bm{\phi}(\bm{x})^{{\rm T}}\bm{\beta}^{[k]*}.$
Let $k(\cdot,\cdot):\mathbb{R}^{p}\times\mathbb{R}^{p}\rightarrow\mathbb{R}$ be the kernel function corresponding to $\bm{\phi}(\bm{x})$, and let $\bm{K}$ be a $p\times p$ semi-definite symmetric matrix whose $(l,m)$-component is $k(x_{l},x_{m})$. If we estimate the parameter $\bm{\alpha}$ that satisfies
$\bm{H}^{[1]}(\bm{d},\bm{x};\bm{\alpha})={\rm E}\bigg[e^{[1]}(\bm{x};\bm{\alpha})\bigg\{\frac{d^{[1]}}{e^{[1]}(\bm{x};\bm{\alpha})}-1\bigg\}\bm{K}\bigg]=\bm{{\rm O}},$
$\bm{H}^{[0]}(\bm{d},\bm{x};\bm{\alpha})={\rm E}\bigg[e^{[1]}(\bm{x};\bm{\alpha})\bigg\{\frac{d^{[0]}}{e^{[0]}(\bm{x};\bm{\alpha})}-1\bigg\}\bm{K}\bigg]=\bm{{\rm O}},$
and then estimate $\bm{\theta}$ and derive a model selection criterion based on it, it would be a natural extension of the proposed method.
There are other valuable directions of study. For example, Callaway and Sant’Anna (2021) and Goodman-Bacon (2021) proposed a DID approach in a more flexible setting than the original one, where there are multiple time points and treatment can be initiated at any of them. Moreover, De Chaisemartin and D’haultfœuille (2023) developed a DID approach for when there are more than two groups, and Richardson et al. (2023) proposed one to estimate causal effects under alternative assumptions to the parallel trend assumption. While the outcome in the basic DID approach is continuous, Tchetgen et al. (2024) dealt with the case where the outcome follows a generalized linear model. In this way, the DID approach has been theoretically developed in accordance with actual problems, so we believe that extending the proposed method to enable it to deal with those problems would thus be a potentially profitable goal for the near future. In econometrics, synthetic control (Abadie et al. 2010) has also become standard in causal inference, and a combination of the two has been developed as synthetic DID (Arkhangelsky et al. 2021). Here, because the objective of the information criterion is what to synthesize and with what weights, the proposed method cannot necessarily be used; however, an investigation along the lines presented in this paper might prove fruitful. Furthermore, propensity score analysis with high-dimensional covariates is also an important topic (e.g., Belloni et al. 2017, Chernozhukov et al. 2018, Ning et al. 2020b), wherein selecting covariates that truly have impact on causal effects and robustly estimating these effects is an essential task to ensure the validity of the analysis. Even when the covariates are high-dimensional, asymptotic properties of the estimator of causal effects tend to be obtained as well, so it is surely possible to extend our method.
Disclosure statement
No potential conflict of interest was reported by the authors.
Data availability statement
The LaLonde dataset is available in the R package Matching version 4.10.14 (Sekhon 2011). Example R scripts used for the simulations and real data analyses are available from the authors.
Funding
Yoshiyuki Ninomiya gratefully acknowledges support from JSPS KAKENHI (grant numbers 23H00809 and 23K18471)